### v1

# Adapting Vision-Language Models for Evaluating World Models

Abstract
World models – generative models that simulate environment dynamics conditioned on past observations and actions – are gaining prominence in planning, simulation, and embodied AI. However, evaluating their rollouts remains a fundamental challenge, requiring fine-grained, temporally grounded assessment of action alignment and semantic consistency – capabilities not captured by existing metrics. Vision-Language Models (VLMs) have shown promise as automatic evaluators of generative content due to their strong multimodal reasoning abilities. Yet, their use in fine-grained, temporally sensitive evaluation tasks remains limited and requires targeted adaptation. We introduce a evaluation protocol targeting two recognition tasks – action recognition and character recognition – each assessed across binary, multiple-choice, and open-ended formats. To support this, we present universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a method for adapting VLMs to rollout evaluation under data and compute constraints. We conduct a large-scale study comparing full, partial, and parameter-efficient finetuning across task formats, context lengths, sampling strategies, and data compositions. The resulting unified evaluator matches the performance of task-specific baselines using a single checkpoint. Human studies confirm strong alignment with human judgments, establishing universe as a scalable, semantics-aware evaluator for world models.
1 Introduction
World models are generative models trained to predict future observations conditioned on past observations and actions [37, 41, 5]. They offer a powerful abstraction for learning, reasoning, and planning in complex interactive environments, and are rapidly becoming foundational in domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77]. As their capabilities grow, a persistent challenge remains: evaluation.
Rollouts produced by world models are visually rich, temporally grounded, and semantically structured. Evaluating them requires: (i) precise alignment between generated frames and control sequences at the timestamp level [106], and (ii) consistent tracking of entities and their semantics over time [56].
Existing evaluation protocols offer limited insight in this regime. Traditional visual generation metrics are ill-suited to the structured and dynamic nature of rollouts: (i) early distributional metrics focus on static images and are sensitive to low-level variations [85, 46, 12], (ii) motion-aware metrics such as FVD [97] capture temporal patterns but lack semantic grounding, and (iii) multimodal metrics improve semantic alignment but ignore action conditioning and timestamp-level fidelity [52]. While human evaluation remains the gold standard [2, 6], it is expensive and difficult to scale. Similarly, emerging text-to-video benchmarks [66, 50, 62] focus on open-ended generation and neglect the fine-grained control and temporal coherence essential to evaluating world model outputs.
[…]
VLM have demonstrated strong generalization across multimodal tasks by integrating visual and linguistic reasoning [61, 30, 24, 99, 1, 64, 26, 72], and are increasingly used to evaluate generative models [59, 70, 63, 20]. We extend this direction, hypothesizing that VLMs can serve as fine-grained evaluators of world model rollouts, capturing semantic consistency, temporal coherence, and alignment with control inputs. Virtual environments offer a favorable testbed for this purpose, as they expose full access to timestamped actions, object states, and agent behaviors, enabling controlled and systematic evaluation.
Figure 1: Performance and efficiency of universe (orange bars throughout) compared to task-specific baselines (multiple colours). Left and Center: Action recognition and Character Recognition accuracy across binary, multiple-choice, and open-ended settings. Right: Sample efficiency – our adaptation recipe achieves strong performance with substantially fewer training samples per epoch. universe matches the performance of task-specific baselines while using a single checkpoint, enabled by mixed supervision, efficient frame sampling, and lightweight partial fine-tuning.
[…]
To support this protocol, we introduce universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a method for adapting VLMs to structured rollout evaluation. universe emerges from a systematic study of adaptation strategies under realistic data and compute constraints. We analyze the impact of supervision regime, frame sampling strategy, visual context length, and training budget, arriving at an adaptation recipe that combines mixed supervision, efficient frame selection, and lightweight fine-tuning.
To assess the reliability of universe, we conduct a human annotation study using rollouts generated by WHAM [56], a publicly available world model trained in a complex interactive environment. We evaluate the rollouts using universe and collect human judgments on the same data to compare responses. We find that universe’s judgments exhibit strong alignment with human ratings, confirming its potential for automated evaluation world model rollouts, particularly when explicit ground truth is unavailable or costly to obtain.
In summary, this work makes three key contributions:
  1. I
    We introduce a structured protocol for evaluating world model rollouts, based on action and character recognition tasks with increasing levels of complexity;
  2. II
[…]
  3. III
    We show that universe matches the performance of task-specific baselines using a single checkpoint, demonstrating its ability to generalize across tasks with a unified model (Figure 1). Furthermore, its predictions align closely with human judgments, validating its effectiveness as a scalable, semantics-aware evaluator for world model rollouts.
2 Related Work
Challenges in Evaluating World Models. World models are generative systems that learn predictive representations of environment dynamics [37], originally proposed for model-based RL [92] and now central to domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77]. Recent models such as Dreamer v1–3 [39, 38, 40, 41], MuZero [86], IRIS [74], UniSim [106], and DIAMOND [5] have improved rollout fidelity and controllability. Yet evaluation largely focuses on downstream success metrics—e.g., game score or goal completion [9, 54, 36, 7, 8]—which provide only coarse, indirect signals of rollout quality. Genie [17, 78] decouples world model learning and agent training, but its evaluation still emphasizes visual quality and control, without probing semantic or causal fidelity. Cosmos [2] proposes a structured protocol that combines FID/FVD with structure-from-motion-based 3D consistency checks and human ratings on instruction following, object permanence, and visual verity. While insightful, this approach is tied to simulator-specific infrastructure and requires costly manual comparison. Human-in-the-loop protocols such as the Video Generation Arena [6] also rely on pairwise comparison to assess rollout quality. These methods, though informative, are expensive and hard to scale.
Evaluation Metrics and Protocols for Visual Generation. Early evaluations of generative models relied on full-reference metrics such as PSNR and SSIM [100], which capture pixel-level and perceptual similarity but are sensitive to spatial misalignments and fail to reflect semantic fidelity. To address this, distributional metrics like Inception Score (IS) [85], Fréchet Inception Distance (FID) [46], and Kernel Inception Distance (KID) [12]. Other proposals such as PPL [57], Parzen likelihoods [34], and HYPE [110] attempt to quantify perceptual smoothness or human realism, but remain focused on static images. For video generation, FVD [97] generalizes FID using I3D features [18], introducing a motion-aware distributional baseline. Yet, FVD also lacks semantic grounding and does not account for causal structure or goal alignment. To improve semantic grounding, metrics based on text-image alignment have been proposed. CLIPScore [45] and CLIPSIM [102] compute similarity between generated visuals and textual or visual references using CLIP embeddings [82], while Jayasumana et al. [52] extend this to distributional comparisons via MMD. However, all operate at the frame level. Structured evaluation protocols using vision-language reasoning have also emerged. VQA Accuracy [70] uses LLMs to score answers on static image questions, and VQAScore [63] probes alignment via templated binary queries. Lee et al. [59] propose VLM evaluator to evaluate other VLMs responses given user criteria. These approaches introduce task structure but remain limited to single-frame evaluation. Recent text-to-video (T2V) benchmarks such as EvalCrafter [66], VBench [50], and DEVIL [62] introduce curated prompts and metrics covering text alignment, motion realism, and perceptual quality. While these protocols push forward evaluation of open-ended video generation, they lack timestamp-level action grounding.
Vision-Language Model Adaptation. VLMs have emerged as powerful tools for multimodal understanding, demonstrating strong performance across tasks such as captioning, retrieval, visual question answering, and instruction following [61, 30, 24, 99, 1, 64, 26, 72]. Adaptation approaches can be broadly categorized into prompt-level and weight-level methods. One prominent prompt-level adaptation techniques is prompt tuning, which injects task information directly into the input space [75, 111, 103], and in-context learning (ICL), where models such as GPT-3 [16] and Flamingo [4] condition on task demonstrations at inference time without updating parameters. Retrieval-augmented generation (RAG) [60] combines parametric models with non-parametric memory, and multimodal variants incorporate external visual or auditory context [49, 22]. While lightweight, these approaches are limited in their ability to model temporal dependencies or align with structured rollouts. Weight-level adaptation enables stronger domain alignment but incurs higher computational cost. Full finetuning remains effective yet costly, while partial finetuning [107] offers a trade-off by updating only selected layers. Parameter-efficient finetuning (PEFT) provides a scalable alternative and can be grouped into low-rank and adapter-based strategies [42]. Low-rank methods, such as LoRA [48], inject rank-constrained updates into frozen layers. Recent extensions improve upon this via weight decomposition [65], quantization-aware adaptation [27, 105], mixture-of-experts routing [104], and long-context support [25]. Adapter-based methods insert lightweight modules between frozen layers to enable modular adaptation with minimal overhead [69, 109]. A parallel line of work investigates multimodal few-shot learning. Frozen [95] was among the first to explore this setting, followed by works combining prompting and ICL for improved sample efficiency [53, 89], and works introducing a learnable meta-mapper to bridge frozen VLM components for few-shot meta-learning [76].
[…]
3 Methodology
We consider the problem of evaluating rollouts generated by world models in interactive environments. A world model $W$ is trained to predict the next observation $o_{t}$ given the past observations $o_{<t}$ and actions $a_{<t}$: $W:(o_{<t},a_{<t})\rightarrow o_{t}$, where $o_{t}\in\mathcal{O}$ represents the sensory observation at timestep $t$, typically an RGB image. Rollouts consist of temporally grounded sequences that reflect the causal effects of control inputs. These outputs are semantically rich and visually complex, requiring timestamp-level assessment of correctness.
To enable automatic evaluation, we propose universe, an adapted Vision-Language Model (VLM) that serves as a structured evaluator for world model rollouts. Formally, it operates as a function: $E:(V,Q)\rightarrow\hat{A}$, where $V=(o_{t_{1}},\dots,o_{t_{k}})\in\mathcal{O}^{k}$ is a sequence of frames from a rollout, $Q\in\mathcal{L}$ is a natural language question, and $\hat{A}\in\mathcal{L}$ is the predicted answer. Evaluation quality is measured by comparing $\hat{A}$ to the reference answer $A$ using semantic similarity metrics.
[…]
Dataset Construction. Effective VLM adaptation for rollout evaluation requires a dataset that (i) captures realistic human behavior in interactive environments, and (ii) aligns with prior work in simulated settings to support comparability and reproducibility. Additionally, since WHAM – the world model used in our evaluation – is trained on Bleeding Edge, it is critical that the adaptation dataset match its distribution. To satisfy these constraints, we partnered with Ninja Theory and curated a dataset from both internal and public Bleeding Edge gameplay recordings, focusing on the Skygarden map used in WHAM [56]. This dataset provides high visual and behavioral diversity [79], includes a publicly available evaluation split, and is closely aligned with prior work in the domain [56, 79, 94, 88, 28], enabling cross-method comparison.
Data preparation proceeds in three stages: (i) Preprocessing:Segment gameplay into 14-frame clips with synchronized video, control logs, and metadata; (ii) Description Generation:Convert structured annotations (e.g., actions, agent states) into natural language summaries; (iii) Question-Answer Pair Construction:Generate six QA pairs per clip (binary, multiple-choice, and open-ended) spanning both AR and CR tasks. The final dataset contains 32.453 training clips and 8.113 validation clips, yielding 194.718 and 48.678 QA pairs, respectively. See Appendix C for details.
Model Architecture. We adapt a model from the PaliGemma family [10, 91], consisting of a vision encoder $\mathcal{M}_{V}$, a projection head $\mathcal{M}_{P}$, and a language decoder $\mathcal{M}_{L}$. Based on initial zero-shot evaluations (Appendix F.1), we use a single configuration for all experiments—PaliGemma 2 3b, which includes a 2B-parameter Gemma 2 decoder pretrained on 2T tokens. Input frames are resized to $224\times 224$ and tokenized into 256 patches each. Model achitecture details are in Appendix D.1.
Each model input sequence $S=\{S_{\mathcal{I}},S_{\mathcal{T}}^{\text{PREF}},S_{\mathcal{T}}^{\text{SUFF}}\}$ consists of: visual tokens $S_{\mathcal{I}}$ from $k$ frames, a textual prefix $S_{\mathcal{T}}^{\text{PREF}}$ containing the task-language cue and question, and a suffix $S_{\mathcal{T}}^{\text{SUFF}}$ with the expected answer (used only during training). This format allows the decoder to attend jointly over visual and textual context. Full prompt details are provided in Appendix D.1.
Training Objective. We optimize a causal language modeling loss on the answer suffix:
$\mathcal{L}(S)=-\sum_{t=1}^{T_{\text{SUFF}}}\log P(s_{t}^{\text{SUFF}}\mid S_{<t^{\prime}})$ (1)
where $s_{t}^{\text{SUFF}}$ is the $t$-th token in the suffix, and $t^{\prime}=T_{\mathcal{I}}+T_{\text{PREF}}+t$ is the token position in the flattened sequence.
Adaptation Strategies. We explore a broad design space for adapting pretrained VLMs to temporally grounded rollout evaluation. Our study spans three core dimensions: fine-tuning configurations, frame sampling policy, and supervision composition.
Fine-Tuning Configurations. We compare five adaptation strategies varying in parameter count and modularity: (i) Zero-shot prompting:No tuning; model is prompted directly. (ii) Full fine-tuning:All parameters $\theta=\theta_{V}\cup\theta_{P}\cup\theta_{L}$ are updated end-to-end. (iii) Dual-component fine-tuning:Two of three modules are trained (e.g., $\theta_{P}\cup\theta_{L}$). (iv) Single-component fine-tuning:Only one module—vision, projection, or language—is updated. (v) Parameter-efficient fine-tuning:We apply LoRA [48] adapters to attention and MLP layers in vision and language components: $\mathbf{W}\leftarrow\mathbf{W}+\frac{\alpha}{r}\mathbf{A}\mathbf{B},\quad\alpha=8,\;r\in\{8,16,32,48,64\}$.
Frame Sampling Policy. We vary both the number of input frames and their sampling strategy. Specifically, we sweep over $k\in[1,8]$, and evaluate two selection methods: (i) First-$n$: selecting the first $k$ frames from each rollout; (ii) Uniform-$n$: sampling $k$ frames uniformly across the full clip.
Supervision Composition. To support generalization across QA formats and tasks, we construct a multi-task dataset covering binary, multiple-choice, and open-ended prompts across both AR and CR. We perform a three-stage grid search to optimize the data mixture: (i) Varying AR/CR task ratios ($\alpha_{\texttt{AR}}$, $\alpha_{\texttt{CR}}$) while fixing QA type proportions ($\beta_{\texttt{Binary}}$, $\beta_{\texttt{MC}}$, $\beta_{\texttt{OE}})$; (ii) Tuning the proportion of open-ended supervision ($\beta_{\texttt{OE}}$) for best performance; (iii) Adjusting $\beta_{\texttt{Binary}}$ and $\beta_{\texttt{MC}}$.
universe: UNIfied Vision-language Evaluator for Rollouts in Simulated Environments. We distill our empirical findings into universe, a lightweight and scalable adaptation method for temporally grounded evaluation of world model rollouts using VLMs. Designed for constrained compute and limited supervision, universe delivers strong generalization across our evaluation protocol using a single, partially tuned model. The method combines three key components:
  1. I
    Partial fine-tuning: We update only the projection head ($\theta_{P}$), training just 0.07% of model parameters. Despite this minimal footprint, it achieves the second-best performance among all strategies—trailing only vision encoder tuning, which requires $\sim$11of parameters and incurs significantly higher compute cost.
  2. II
    Efficient frame sampling: Each input sequence includes $k=8$ frames sampled uniformly from a 14-frame rollout. This sparsity-aware strategy maintains long-range temporal structure while reducing token count and enabling efficient batching.
  3. III
    Mixed supervision: We train on a hierarchical mixture of tasks and QA formats. The task distribution favors Action Recognition ($\alpha_{\texttt{AR}}=0.8$) due to its stronger causal grounding. Within each task, we emphasize open-ended questions ($\beta_{\texttt{OE}}=0.8$), while maintaining smaller proportions of binary ($\beta_{\texttt{binary}}=0.15$) and multiple-choice ($\beta_{\texttt{MC}}=0.05$) examples.
4 Experiments
[…]
Figure 2: Comparison of universe and baseline models on Action and Character Recognition. Left: universe outperforms all baselines on AR. Right: On CR, it ranks third, behind models with either full vision encoder tuning or task-specific training with greater supervision. Trained under a unified protocol with minimal parameter updates (0.07%) and reduced per-task data, universe delivers strong performance across both tasks, highlighting its efficiency and generalization.
Baselines. We compare universe against two classes of baselines: (i) Zero-shot VLMs:Seven off-the-shelf models, including VideoLLaMA3 (2B, 7B) [14] and three PaliGemma models: version 1 (3B) and version 2 (3B and 10B) [10, 91], evaluated without domain adaptation using an 8-frame visual context window. [^1] We also experimented with CLIPScore-based evaluation (Appendix F.2); results underperformed relative to selected baselines and were constrained to predefined candidate sets, further underscoring the need for model adaptation. (ii) Fine-tuned PaliGemma 2:Variants adapted via full, partial, and parameter-efficient tuning. This backbone is selected based on a sweep over PaliGemma variants, using zero-shot performance as a guide (Appendix F.1). The adaptation space includes 8 primary baselines: (i) Single-component fine-tuning: tuning only the vision encoder ($\mathcal{F}_{\text{V}}$), the multimodal projector ($\mathcal{F}_{\text{P}}$), or the language head ($\mathcal{F}_{\text{L}}$); (ii) Two-component fine-tuning: jointly tuning pairs of components—$\mathcal{F}_{\text{V+P}}$, $\mathcal{F}_{\text{V+L}}$, and $\mathcal{F}_{\text{P+L}}$; (iii) Full-model fine-tuning: tuning all components simultaneously ($\mathcal{F}_{\text{V+P+L}}$); (iv) LoRA-based tuning: Parameter-efficient adaptation with rank $r=8$, selected after observing minimal performance variation across $r\in\{8,16,32,48,64\}$ (see Appendix F.3 for details). All models are trained using 8-frame clips and a single epoch.
Results. Figure 2 (left, center) summarizes performance across Action Recognition (AR) and Character Recognition (CR). Zero-shot models perform poorly across both tasks: VideoLLaMA3 variants fall below 12.7% on AR and 6.4% on CR; PaliGemma variants reach up to 29.7% on AR and 17.2% on CR. These results confirm that general-purpose VLMs lack the temporal grounding and domain-specific semantics required for structured rollout evaluation. In contrast, universe achieves strong performance—outperforming all models on AR and ranking third on CR. The top two CR models either fine-tune the full vision encoder, requiring updates to $\sim$400M parameters, or fine-tune the multimodal projector with 5 $\times$ more CR supervision, each trained in isolation on a single task and prompt format. universe, by comparison, tunes only the 2.66M-parameter projector (0.07% of the model), under a unified protocol spanning both tasks, all prompt formats, and reduced per-task supervision. Its performance under these constraints underscores the efficiency and generality of our adaptation strategy for temporally grounded evaluation.
5 Analysis
[…]
Supervision and Temporal Context. We begin by analyzing how supervision (training budget) and temporal input (number of frames) influence universe performance. By independently and jointly varying the number of training epochs and input frames, we disentangle the contributions of model capacity and temporal context to task success.
Figure 3: Exact Match accuracy over training epochs for Action Recognition (AR, top) and Character Recognition (CR, bottom). AR improves gradually with supervision, while CR converges to high accuracy with minimal training.
[…]
Figure 4: Action Recognition performance as a function of training supervision (epochs) and temporal context (number of frames), evaluated across all formats. Performance improves along both axes, with highest accuracy achieved when both dimensions are scaled.
[…]
Results. As shown in Figure 5, uniform-$n$ consistently outperforms first-$n$ across all evaluation formats. The effect is most pronounced at low frame counts. With only 2 input frames, uniform sampling improves exact match accuracy from 84.42% to 90.47% in Binary, from 65.53% to 83.93% in Multiple-Choice, and from 65.38% to 82.68% in Open-Ended formats. Gains persist even at 8 frames, where uniform sampling maintains an advantage across formats.
Figure 5: Effect of frame sampling strategy on Action Recognition performance across all formats. Uniform-$n$ sampling (orange) consistently outperforms first-$n$ (blue), with especially large gains at low frame counts, and maintains an advantage as temporal context increases.
[…]
Figure 6: Hierarchical ablation of training data composition for universe. Left: Varying task-level ratio $\alpha$ (AR vs. CR) with uniform format distribution ($\beta=1/3$) shows that increasing $\alpha_{\texttt{AR}}$ improves AR performance, especially for multiple-choice, while open-ended remains flat. Center: Sweeping format-level ratio $\beta_{\texttt{OE}}$ with fixed $\alpha_{\texttt{AR}}=0.8$ reveals that oversampling open-ended data ($\beta_{\texttt{OE}}=0.8$) improves AR-OE performance. Right: Fine-tuning binary and MC proportions under $\beta_{\texttt{OE}}=0.8$ shows performance is stable across mixes, with slight gains from $\beta_{\texttt{binary}}=0.15$, $\beta_{\texttt{MC}}=0.05$.
Figure 7: Comparison of two training regimes: a default data mix (equal task and format proportions) and the optimized mix derived from hierarchical tuning. The optimized configuration yields substantial gains on AR, while maintaining strong CR performance.
Data Mix Optimization. To determine an effective training mixture for universe, we perform a hierarchical ablation over task-level and format-level data ratios. We begin by varying the task-level proportion $\alpha$ (AR vs. CR), holding the format distribution fixed at $\beta_{\texttt{binary}}=\beta_{\texttt{MC}}=\beta_{\texttt{OE}}=\nicefrac{{1}}{{3}}$. As shown in Figure 6 (left), increasing $\alpha_{\texttt{AR}}$ improves AR performance—especially for multiple-choice—while CR remains stable, with a favorable tradeoff reached at $\alpha_{\texttt{AR}}=0.8$. However, open-ended accuracy shows little change, motivating format-specific rebalancing. Fixing $\alpha_{\texttt{AR}}=0.8$, we sweep the format ratio $\beta_{\texttt{OE}}$, and observe in Figure 6 (center) that AR-OE accuracy improves substantially with increased open-ended coverage, peaking at $\beta_{\texttt{OE}}=0.8$, albeit at the cost of binary performance. To restore balance, we fix $\beta_{\texttt{OE}}=0.8$ and allocate the remaining budget across binary and multiple-choice formats. As shown in Figure 6 (right), performance remains robust across configurations, with a slight preference for $\beta_{\texttt{binary}}=0.15$ and $\beta_{\texttt{MC}}=0.05$. Based on these findings, we adopt the following optimized data composition: $\alpha_{\texttt{AR}}=0.8$, $\alpha_{\texttt{CR}}=0.2$; $\beta_{\texttt{binary}}=0.15$, $\beta_{\texttt{MC}}=0.05$, and $\beta_{\texttt{OE}}=0.8$.
Effectiveness of the Optimized Mix. Having identified an optimized training mixture through hierarchical ablation, we now evaluate its impact in practice. We compare the final universe model—trained with this optimized mix—to a baseline trained with a default task and format distribution. We train both models on 4 epochs. As shown in Figure 7, the optimized configuration yields substantial gains on AR, particularly for multiple-choice and open-ended formats, while maintaining competitive performance on CR. These results underscore the importance of data composition in enabling robust multi-task learning within a unified evaluator.
6 Evaluating World Model Rollouts with universe
We evaluate universe as an automated evaluator of world model rollouts using the WHAM benchmark [56], which provides pretrained world models and an evaluation set. Our analysis focuses on two axes: (i) in-domain accuracy, measured on Skygarden—the environment used during fine-tuning—and (ii) generalization to six unseen environments.
We compare universe’s predictions on samples from two world models: (i) WHAM-140M, trained on Skygarden with lower-quality rollouts ($128\times 128$ resolution), and (ii) WHAM-1.6B, trained on a diverse environment suite with higher-resolution output ($300\times 180$). We prepare 30 rollouts for each model-environment pair, yielding a total of 240 rollouts. Each rollout is segmented into 14-frame clips and paired with six natural language questions, following our evaluation protocol. universe answers each question using majority voting over five greedy decoding samples. Human annotators then rate each response on a four-point ordinal scale: Correct, Partially Correct, Incorrect, and Unclear. Each response is independently rated by two annotators; in cases of disagreement, a third annotator serves as adjudicator. Inter-annotator agreement is quantified using Cohen’s $\kappa$. Full details of the annotation protocol are provided in Appendix E.
Results. Figure 8 summarizes graded accuracy across models and environments. We observe a significant performance gap between rollouts from WHAM-140M and WHAM-1.6B. Despite being in-domain, WHAM-140M yields lower evaluation accuracy likley due to a resolution mismatch: its $128\times 128$ frames are upsampled to $224\times 224$ for universe input. In contrast, WHAM-1.6B produces higher-quality inputs, enabling more reliable evaluation. On WHAM-1.6B Skygarden rollouts, universe achieves 75.02% graded accuracy on AR and 90.00% on CR. When applied to six previously unseen environments, AR accuracy remains relatively stable with the lowest accuracy of 66.11% for environment G, and highest accuracy of 83.33% for environment E. Overall, the results indicate strong generalization, especially for identity-focused recognition tasks. Cohen’s $\kappa$ scores further reflect the interpretability of universe’s outputs: agreement is substantial overall ($\kappa=0.73$), with the lowest in Environment G ($\kappa=0.59$) and highest in Environment A on WHAM-1.6B rollouts ($\kappa=0.91$). These results highlight that universe remains aligned with human judgments.
Figure 8: Graded accuracy of universe across rollouts from WHAM-140M (Env. A) and WHAM-1.6B (Envs. A–G). Performance improves markedly with higher-fidelity rollouts from WHAM-1.6B, even in out-of-domain settings (B–G). Cohen’s $\kappa$ (above bars) reflects inter-rater agreement.
7 Conclusion
In this paper, we investigate the use of Vision-Language Models (VLMs) as automated evaluators for world model rollouts, addressing the fundamental challenge of fine-grained, temporally grounded evaluation. We introduce a structured evaluation protocol centered on action and character recognition tasks across binary, multiple-choice, and open-ended formats. To support this, we propose universe, a unified method for adapting VLMs to this setting through mixed supervision, efficient frame sampling, and lightweight fine-tuning. Our large-scale study demonstrates that universe matches the performance of task-specific baselines using a single checkpoint and aligns closely with human judgments, establishing it as a scalable, semantics-aware evaluator for evaluating world models, particularly when explicit ground truth is unavailable or costly to obtain.
Limitations While universe performs well in simulation, its generalization beyond simulated environments remains an open challenge. The evaluation protocol targets two fidelity axes, which, while comprehensive, omit higher-order reasoning over goals, causality, and multi-agent dynamics. Our experiments focus on short to medium context lengths; scaling to long-horizon rollouts remains an open challenge, especially under limited supervision. Although compute-efficient, training could be further improved with adaptive curricula or progressive tuning. Finally, like all pretrained VLMs, universe may reflect dataset biases and underperform on rare or ambiguous behaviors.
Supplementary Materials
[…]
Appendix A Broader Impact
[…]
This capability has direct implications for high-impact domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77], where world models simulate environment dynamics and support downstream control and generalization. In such contexts, precise and interpretable evaluation is critical not only for benchmarking, but also for diagnosing failure modes and ensuring alignment with intended behaviors.
[…]
We therefore advocate for cautious deployment, accompanied by human oversight, rigorous validation, and transparent reporting. While universe advances the automation of world model evaluation, it must be situated within evaluation pipelines that foreground robustness, interpretability, and accountability.
Appendix B Reproducibility Statement
To support reproducibility and facilitate future research, we provide detailed instructions for reproducing all main experiments. Detailed descriptions of model architectures, training procedures, and dataset construction are provided in Section 4 and Appendix D. All experiments have been repeated for three runs. Plots and tables with quantitative results show the standard deviation across these runs.
[…]
Compute Resources. All experiments were conducted using NVIDIA A100 GPUs (40GB memory) on an internal compute cluster. Each model was trained and/or evaluated using 8 GPUs. The compute breakdown is as follows: zero-shot evaluation experiments consumed approximately 136 GPU-days; baseline fine-tuning experiments required around 864 GPU-days; analysis experiments contributed the bulk of usage, totaling 2,554 GPU-days. Human evaluation experiments—including rollout generation and response annotation using universe—incurred an additional 1.125 GPU-days. Additional compute was required for preliminary experiments, and failed runs not included in the final paper. These development activities accounted for an estimated 1,599 GPU-days. In total, all experiments amounted to approximately 5,153 GPU-days, equivalent to 14.12 GPU-years.
Appendix C Dataset
[…]
C.1 Construction Process
The ground truth dataset for adapting the evaluator (see Section 3) was developed in collaboration with Ninja Theory using human gameplay recordings from Bleeding Edge, a 4v4 multiplayer combat game. Data use was governed by a formal agreement with the studio, and collection adhered to the game’s End User License Agreement (EULA). All protocols were approved by our Institutional Review Board (IRB), and personally identifiable information (PII) was removed prior to analysis.
Each gameplay session is represented as a tuple $s=(v,c,m)$, where $v$ is a high-resolution MP4 video (60 FPS), $c$ is the synchronized controller action log, and $m$ contains structured metadata (e.g., player roles, agent identities, action categories, and map context). The full set of gameplay sessions is denoted by $\mathcal{S}=\{(v_{i},c_{i},m_{i})\}_{i=1}^{|\mathcal{S}|}$.
The dataset construction pipeline proceeds in three stages:
  1. (i)
    Preprocessing. We begin by filtering out corrupted applying or inactive sessions and synchronizes the video, controller logs, and metadata streams using internal game timestamps: $\mathcal{S}_{\text{valid}}=\texttt{Preprocessing}(\mathcal{S})$. Each valid session is segmented into non-overlapping clips of fixed length $L=14$ frames, each paired with controller input and shared metadata; formally, for a session $s=(v,c,m)\in\mathcal{S}_{\text{valid}}$, the segmentation produces $\texttt{Segment}(v,c,m,L)=\{(f^{(1:L)},c^{(1:L)},m)\}$, where $f^{(1:L)}$ denotes the sequence of frames, $c^{(1:L)}$ the aligned controller inputs, and $m$ the associated metadata. The complete set of extracted clips across all valid sessions is defined as $\mathcal{V}=\bigcup_{s\in\mathcal{S}_{\text{valid}}}\texttt{Segment}(s,L)$, where each element $v\in\mathcal{V}$ is a triplet $(f^{(1:L)},c^{(1:L)},m)$ consisting of video frames, corresponding controller inputs, and metadata.
  2. (ii)
    Description Generation. Next, for each sequence of frames $f^{(1:L)}\in\mathcal{V}$, we use the associated control log $c^{(1:L)}$ to extract action information and the metadata $m$ to obtain character-related attributes. These are combined to generate a structured natural language description via $d=\texttt{Describe}(c^{(1:L)},m)$. This yields a set of paired video–text examples: $\mathcal{Z}=\left\{(f^{(1:L)},d)\mid f^{(1:L)}\in\mathcal{V}\right\}$.
  3. (iii)
    Question-Answer Pair Construction. Finally, we generate six QA pairs per clip, spanning two predefined tasks (AR and CR), each instantiated in three question formats: binary, multiple-choice, and open-ended. To enable this, we define task-specific answer spaces using $\texttt{GetAnswerSpace}(\mathcal{Z})$, which returns $\mathcal{Y}_{\text{AR}}$ for action categories and $\mathcal{Y}_{\text{CR}}$ for character identities, based on all video–text pairs in $\mathcal{Z}$. For each clip, we extract the task-specific ground-truth answer from the corresponding description as $y=\texttt{ExtractLabel}(d,t)$, where $t\in\{\text{AR},\text{CR}\}$. Each QA format is constructed as follows: (i) Binary: Two binary question-answer pairs are generated per instance using FormatBinaryPrompt. The positive question $Q^{\text{pos}}$ is constructed using the correct label $y\in\mathcal{Y}^{(t)}$ and paired with the positive answer $A^{\text{pos}}$. The negative question $Q^{\text{neg}}$ is constructed using an incorrect label $\tilde{y}\sim\texttt{SampleDistractor}(\mathcal{Y}^{(t)}\setminus\{y\})$ and paired with the negative answer $A^{\text{neg}}$. (ii) Multiple-Choice: A question $Q$ is generated using the full set of candidate options, formatted via FormatOptions$(\mathcal{Y}_{t})$. The question is constructed with FormatMCPrompt$(t,O)$ and paired with the correct answer $y\in\mathcal{Y}_{t}$. (iii) Open-Ended: A free-form question $Q$ is generated using FormatOEPrompt$(t)$, prompting the model to produce the correct label $y\in\mathcal{Y}_{t}$ without access to predefined answer choices.
The final dataset is represented as $\mathcal{D}=\{(f_{i}^{(1:L)},QA_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $f^{(1:L)}$ is a video clip and $QA=\{(Q_{j},A_{j})\}_{j=1}^{6}$ is the associated set of question–answer pairs, covering all combinations of three question formats (binary, multiple-choice, open-ended) and two tasks (Action Recognition and Character Recognition). A detailed data pipeline is provided in Algorithm 1.
Procedure DatasetCreation($\mathcal{S},L$):
       $\mathcal{S}_{\text{valid}}\leftarrow\texttt{Preprocessing}(\mathcal{S})$
      $\mathcal{V}\leftarrow\emptyset$
      for $(v,c,m)\in\mathcal{S}_{\text{valid}}$ do
             $\mathcal{V}_{s}\leftarrow\texttt{Segment}(v,c,m,L)$
            $\mathcal{V}\leftarrow\mathcal{V}\cup\mathcal{V}_{s}$
      $\mathcal{Z}\leftarrow\emptyset$
      for $(f^{(1:L)},c^{(1:L)},m)\in\mathcal{V}$ do
             $d\leftarrow\texttt{Describe}(m,c^{(1:L)})$
            $\mathcal{Z}\leftarrow\mathcal{Z}\cup\{(f^{(1:L)},d)\}$
      $\mathcal{D}\leftarrow\emptyset$
      $\mathcal{Y}_{\text{AR}},\mathcal{Y}_{\text{CR}}\leftarrow\texttt{GetAnswerSpace}(\mathcal{Z})$
      for $(f^{(1:L)},d)\in\mathcal{V}$ do
             $\mathcal{QA}\leftarrow\texttt{GenerateQAPairs}(d,\mathcal{Y}_{\text{AR}},\mathcal{Y}_{\text{CR}})$
            for $(Q,A)\in\mathcal{QA}$ do
                   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(f^{(1:L)},Q,A)\}$
      return $\mathcal{D}$
Procedure GenerateQAPairs($d,\mathcal{Y}_{\text{AR}},\mathcal{Y}_{\text{CR}}$):
       $\mathcal{QA}\leftarrow\emptyset$
      for $t\in\{\text{AR},\text{CR}\}$ do
             $y\leftarrow\texttt{ExtractLabel}(d,t)$
            $QA_{\text{bin}}^{pos},QA_{\text{bin}}^{neg}\leftarrow\texttt{CreateBinaryQA}(t,y)$
            $\mathcal{QA}\leftarrow\mathcal{QA}\cup\{QA_{\text{bin}}^{\text{pos}},QA_{\text{bin}}^{\text{neg}}\}$
            $QA_{\text{mc}}\leftarrow\texttt{CreateMCQA}(t,y,\mathcal{Y}_{t})$
            $\mathcal{QA}\leftarrow\mathcal{QA}\cup QA_{\text{mc}}$
            $QA_{\text{oe}}\leftarrow\texttt{CreateOpenEndedQA}(t,y)$
            $\mathcal{QA}\leftarrow\mathcal{QA}\cup QA_{\text{oe}}$
      return $\mathcal{Q}$
Procedure CreateBinaryQA($t,y$):
       $\tilde{y}\leftarrow\texttt{SampleDistractor}(\mathcal{Y}_{t}\setminus\{y\})$
      $Q^{pos}\leftarrow\texttt{FormatBinaryPrompt}(t,y)$
      $Q^{neg}\leftarrow\texttt{FormatBinaryPrompt}(t,\tilde{y})$
      return $\{(Q^{pos},A^{pos}),\;(Q_{\tilde{y}},A^{neg})\}$
Procedure CreateMCQA($t,y,\mathcal{Y}_{t}$):
       $O\leftarrow\texttt{FormatOptions}(\mathcal{Y}_{t})$
      $Q\leftarrow\texttt{FormatMCPrompt}(t,O)$
      return $Q,y$
Procedure CreateOpenEndedQA($t,y$):
       $Q\leftarrow\texttt{FormatOEPrompt}(t)$
      return $Q,y$
Algorithm 1 Dataset Construction Process
Appendix D Experimental Details
In this section, we provide a detailed description of the dataset preparation process, model architecture, prompt templates, training procedure. Additionally, we provide an overview of all results presented in the main paper in numerical table form, an report additional experimental results leveraging alternate fine-tuning solutions.
D.1 Model
This section provides extended details on the architecture, pretraining configuration, and input formatting of the vision-language models used in our experiments. Our primary backbone is PaliGemma [10, 91].
Table 1: Detailed architecture of the PaliGemma model, comprising a SigLIP-So400m vision tower, a multimodal projection head, and a Gemma-based language decoder. All transformer layers follow standard design and include residual connections around attention and MLP blocks.
Component Configuration
Vision Tower: SigLIP-So400m
Patch Embedding Conv2d(in=3, out=1152, kernel=14, stride=14)
Position Embedding Embedding(num_embeddings=256, emb_dim=1152)
Encoder 27 × Transformer Encoder Layers
   Self-Attention
    Query / Key / Value projection Linear(1152 $\rightarrow$ 1152, bias=True)
   Layer Normalization LayerNorm((1152,), eps=1e-6)
   MLP Block
    Activation Function GELU-Tanh
    Feedforward layer (up) Linear(1152 $\rightarrow$ 4304, bias=True)
    Feedforward layer (down) Linear(4304 $\rightarrow$ 1152, bias=True)
   Layer Normalization LayerNorm((1152,), eps=1e-6)
Post-Encoder Layer Norm LayerNorm((1152,), eps=1e-6)
Multimodal Projection Head
Linear Projection Linear(1152 $\rightarrow$ 2304, bias=True)
Language Model: Gemma
Token Embedding Embedding(vocab=257216, dim=2304)
Decoder Stack 26 × Transformer Decoder Layers
   Self-Attention
    Query projection Linear(2304 $\rightarrow$ 2048, bias=False)
    Key projection Linear(2304 $\rightarrow$ 1024, bias=False)
    Value projection Linear(2304 $\rightarrow$ 1024, bias=False)
    Output projection Linear(2048 $\rightarrow$ 2304, bias=False)
   MLP Block
    Gating projection Linear(2304 $\rightarrow$ 9216, bias=True)
    Down projection Linear(2304 $\rightarrow$ 9216, bias=True)
    Up projection Linear(9216 $\rightarrow$ 2304, bias=True)
    Activation Function GELU-Tanh
   Normalization Layers
    Input Norm RMSNorm(2304, eps=1e-6)
    Post-Attn Norm RMSNorm(2304, eps=1e-6)
    Pre-FFN Norm RMSNorm(2304, eps=1e-6)
    Post-FFN Norm RMSNorm(2304, eps=1e-6)
Rotary Embeddings GemmaRotaryEmbedding
LM Head Linear(2304 $\rightarrow$ 257216, bias=False)
D.1.1 Overview
[…]
Vision Encoder: SigLIP-So400m. The visual backbone $\mathcal{M}_{V}$ is a ViT-style encoder pretrained using a Sigmoid contrastive loss (SigLIP). It processes input images by dividing them into non-overlapping 14 $\times$ 14 patches. Each patch is linearly projected into a 1152-dimensional embedding via a convolutional stem. To encode spatial structure, learned positional embeddings are added before the representation is passed through a stack of 27 SigLIP encoder layers. Each encoder layer contains multi-head self-attention with projection layers for queries, keys, and values, followed by an MLP block with GELU-Tanh activations. All transformer blocks use LayerNorm and residual connections. The vision tower supports multiple input resolutions (224, 448, 896), though our experiments fix resolution at 224px2 for consistency and efficiency.
Multimodal Projection Head. The projection head $\mathcal{M}_{P}$ is a lightweight linear mapping from the vision encoder’s output dimension (1152) to the language decoder’s input dimension (2304). It contains approximately 2.66M parameters and is initialized with zero-mean weights. This head enables alignment between visual and linguistic modalities and is important for bridging the representation gap between the vision and language components.
Language Decoder: Gemma. The language module $\mathcal{M}_{L}$ is a decoder-only Transformer with 26 layers and 2304-dimensional hidden states. Token embeddings are learned over a vocabulary of 257,216 tokens, encoded using the SentencePiece tokenizer [58]. Each Transformer block contains a self-attention mechanism with separate linear projections for queries, keys, and values. The MLP block follows a gated architecture, where the input is processed through parallel down projection and gating projection layers, modulated by a GELU-Tanh activation [44], combined via elementwise multiplication, and then passed through an up projection to return to the model’s hidden dimension. RMSNorm is applied before and after both attention and MLP sublayers to stabilize training. Rotary positional embeddings are added to enable relative position encoding. Output tokens are produced via a tied language modeling head that projects back to the vocabulary space.
D.1.2 Configurations
Table 2 summarizes the architecture components and parameter counts of the PaliGemma configurations available for experimentation. While we focus on the PaliGemma 2 3b variant in our study, we include all publicly released configurations for completeness and to clarify how our selected model compares to other available options. All three variants share the same vision encoder and multimodal integration strategy, differing only in the language decoder. The first configuration, PaliGemma 1 3b, pairs the visual encoder with Gemma 1 (2B), pretrained on 6 trillion tokens, resulting in a total model size of approximately 3 billion parameters. The second configuration, PaliGemma 2 3b, replaces the decoder with Gemma 2 (2B), pretrained on 2 trillion tokens, and maintains a comparable total parameter count. The third and largest variant, PaliGemma 2 10b, uses Gemma 2 (9B) as the decoder, pretrained on 8 trillion tokens, yielding a total model size of approximately 9.7 billion parameters.
Table 2: Component-wise parameter overview of the PaliGemma model.
Component Model / Variant Details # Params
Vision Encoder SigLIP-So400m Input resolutions: 224px2, 448px2, 896px2 400M
Multimodal Projection Connects vision and language components 2.66M
Language Model PG 1 Gemma 1 2B, pre-trained on 6T tokens 3B
PG 2 Gemma 2 2B, pre-trained on 2T tokens 3B
PG 3 Gemma 2 9B, pre-trained on 8T tokens 9.7B
D.1.3 Prompt Format
To generate textual responses, we adopt a unified prompt format for the decoder. Each input sequence consists of image tokens $S_{\mathcal{I}}$, a textual prefix $S_{\mathcal{T}}^{\text{PREF}}$ containing the question, and a suffix $S_{\mathcal{T}}^{\text{SUFF}}$ containing the expected answer. The model autoregressively generates the answer tokens, and training loss is applied only to the suffix.
Let $n$ denote the number of input frames and $p$ the number of visual tokens (patch embeddings) per frame. In our setting, each frame is encoded as $p=256$ visual tokens. The overall input schema is as follows:
$S=\;$ $\underbrace{\texttt{<image>}_{1}^{(1)},\ldots,\texttt{<image>}_{p}^{(1)},\;\ldots,\;\texttt{<image>}_{1}^{(n)},\ldots,\texttt{<image>}_{p}^{(n)}}_{S_{\mathcal{I}}\text{: Visual tokens from }n\text{ frames, each represented as }p\text{ patches}}$
$\underbrace{\texttt{<BOS>},\;\texttt{answer en},\;\texttt{<QUESTION>},\;\texttt{<SEP>}}_{S_{\mathcal{T}}^{\text{PREF}}\text{: Prefix (cue+ question)}}$
$\underbrace{\texttt{<ANSWER>},\;\texttt{<EOS>},\;\texttt{<PAD>},\ldots,\texttt{<PAD>}}_{S_{\mathcal{T}}^{\text{SUFF}}\text{: Suffix (answer)}}$
Here, $S_{\mathcal{I}}$ contains visual tokens produced by the vision encoder $\mathcal{M}_{V}$, and projected into $\mathcal{M}_{L}$ space using $\mathcal{M}_{P}$. The prefix $S_{\mathcal{T}}^{\text{PREF}}$ starts with a special <BOS> token and includes a task-language cue (e.g., ‘‘answer en’’), the question, and a separator <SEP>. The suffix $S_{\mathcal{T}}^{\text{SUFF}}$ contains the target answer, terminated with <EOS> and padded with <PAD> tokens for batching.
D.1.4 Pretraining Data and Filtering
PaliGemma is pretrained on a mixture of large-scale vision-language datasets, including WebLI [23], CC3M-35L [87], VQ2A-CC3M-35L [19], OpenImages [80], and WIT [90]. Data quality and safety are maintained through pornographic content filtering, text safety and toxicity filtering, and privacy-preserving measures.
D.2 Evaluation Metrics
In this section, we provide additional details on metrics used for quantitative evaluation. We employ two complementary metrics: Exact Match (EM) and ROUGE-F1 (ROUGE), which together capture both syntactic precision and semantic alignment.
Exact Match Accuracy ($EM$) measures whether the generated answer is identical to the expected answer, providing a high-precision signal for correctness. Formally, it is defined as:
$EM=\mathbbm{1}{}(\hat{A}=A)$ (2)
where $\hat{A}$ is the model’s prediction and $A$ is the corresponding ground-truth answer. This metric is especially informative for binary and multiple-choice formats where the output space is well-defined.
ROUGE F1 (ROUGE) captures token-level semantic overlap between generated and reference responses by computing the harmonic mean of precision and recall. This allows us to account for partially correct or paraphrased answers. For binary questions, we compute the metric on the bigram level, while for multiple-choice and open-ended formats, we use trigram-level evaluation.
Formally, let $G$ and $R$ denote the sets of $n$-grams in the generated and reference answers, respectively. Precision and recall are defined as:
$P=\frac{|G\cap R|}{|G|},\quad R=\frac{|G\cap R|}{|R|}$ (3)
where $|G\cap R|$ counts overlapping $n$-grams. The ROUGE score is then computed as:
$\text{ROUGE}=2\times\frac{P\times R}{P+R}$ (4)
Together, these metrics provide a robust view of model performance: EM reflects exact correctness, while ROUGE provides a softer measure of semantic fidelity, particularly useful for evaluating open-ended generations.
D.3 Results
Table 3: Summary of hyperparameters used in our experiments.
Hyperparameter Value
Input resolution $224\times 224$
Image frames per input 1–8
Number of epochs 1–10
Batch size (per device) 1
Gradient accumulation steps 4
Optimizer AdamW [67]
Learning rate $5\times 10^{-5}$, cosine annealing
Learning rate warmup 10%
Weight decay $1\times 10^{-6}$
Gradient clipping Global norm, threshold 1.0
VLM backbone PaliGemma 2 (3B) [10]
Hyperparameters. Table 3 summarizes the core training hyperparameters used across all adaptation experiments. We train all models on 8 NVIDIA A100 GPUs with a batch size of 1 per device and accumulate gradients over 4 steps, yielding an effective batch size of 32. Each epoch corresponds to a full pass over the adaptation dataset, and no early stopping is applied. Models were trained for 1–10 epochs depending on task and setting. Optimization is performed using AdamW [67] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, a base learning rate of $5\times 10^{-5}$, and weight decay of $1\times 10^{-6}$. We use cosine learning rate annealing [68] with a linear warmup over the first 10% of training steps. To stabilize training, we apply gradient clipping with a global norm threshold of 1.0. All models use PaliGemma 2 (3B) [10] as the vision-language backbone unless otherwise noted. We vary the number of input frames between 1 and 8 depending on task, and all images are resized to a fixed resolution of $224\times 224$. Training is conducted in bfloat16 precision using data parallelism. Model selection is based on final validation accuracy.
Tabular Results Summary. The following tables summarize primary experimental findings across our study. Each entry corresponds to a core evaluation or analysis in the paper, organized by experimental section and aligned with the corresponding table description.
  • Zero-Shot Evaluation (Section 4): Table 4 reports ROUGE-F1 zero-shot performance of pretrained PaliGemma and VideoLLaMA3 models on Action and Character Recognition tasks. Models are evaluated in a zero-shot setting with 1 or 8 input frames, across binary, multiple-choice, and open-ended formats.
  • Fine-Tuned Baselines (Section 4): Table 5 reports ROUGE-F1 and Exact Match performance of PaliGemma 2 variants fine-tuned using full, partial, and parameter-efficient strategies. All models are trained on a single frame for one epoch, and evaluated across binary, multiple-choice, and open-ended formats.
  • Analysis: Supervision and Temporal Context (Section 5): Table 6 examines early-stage learning dynamics on Character Recognition (CR), with evaluation at sub-epoch intervals. Table 7 reports AR performance as a function of training budget, scaling the number of epochs with a single input frame. Table 8 extends this analysis to jointly vary training epochs and the number of input frames, disentangling the effects of temporal context and supervision on AR.
  • Analysis: Temporal Sampling Strategies (Section 4): Table 9 compares first-$n$ and uniform-$n$ frame sampling strategies for Action Recognition, evaluating model performance across varying temporal context lengths ($n\in\{1,\dots,8\}$).
  • Analysis: Optimizing Data Mix for Unified Multi-Task Evaluation (Section 5): This analysis spans three tables. Table 10 explores task-level trade-offs when jointly training on Action and Character Recognition by varying $\alpha_{\texttt{AR}}$ vs. $\alpha_{\texttt{CR}}$, with format distribution held uniform. Table 11 fixes $\alpha_{\texttt{AR}}=0.8$ and searches over format-level ratios ($\beta$), revealing the impact of increased open-ended (OE) supervision. Table 12 further investigates this high-OE regime, balancing the remaining budget between binary and multiple-choice for optimal performance.
Table 4: Zero-shot ROUGE-F1-based evaluation of PaliGemma (PG) and VideoLLaMA3 (VL3) models on Action and Character Recognition tasks using 1 and 8 input frames. “MC” denotes multiple-choice and “OE” open-ended formats.
Action Recognition Character Recognition
Fr Model Binary MC OE Binary MC OE
1 PG 1 3B 50.43 $\pm$ 0.13 08.12 $\pm$ 0.02 10.83 $\pm$ 0.01 50.73 $\pm$ 0.38 00.46 $\pm$ 0.06 $0$$\pm$ 0.00
PG 2 3B 44.69 $\pm$ 0.03 09.30 $\pm$ 0.17 12.64 $\pm$ 0.01 48.58 $\pm$ 0.07 00.28 $\pm$ 0.06 00.01 $\pm$ 0.00
PG 2 10B 50.04 $\pm$ 0.03 26.98 $\pm$ 0.00 12.35 $\pm$ 0.21 50.08 $\pm$ 0.07 08.33 $\pm$ 0.50 $0$$\pm$ 0.00
VL3-2B 03.24 $\pm$ 0.00 18.52 $\pm$ 0.06 06.27 $\pm$ 0.05 08.76 $\pm$ 0.04 03.44 $\pm$ 0.08 00.50 $\pm$ 0.01
VL3-7B 45.02 $\pm$ 0.28 15.53 $\pm$ 0.05 06.54 $\pm$ 0.04 39.09 $\pm$ 0.73 06.21 $\pm$ 0.05 00.51 $\pm$ 0.02
8 PG 1 3B 51.67 $\pm$ 0.02 10.68 $\pm$ 0.00 10.32 $\pm$ 0.00 51.39 $\pm$ 0.07 00.25 $\pm$ 0.00 $0$$\pm$ 0.00
PG 2 3B 47.61 $\pm$ 0.19 06.73 $\pm$ 0.04 14.52 $\pm$ 0.00 48.37 $\pm$ 0.18 00.03 $\pm$ 0.00 00.01 $\pm$ 0.00
PG 2 10B 50.02 $\pm$ 0.06 26.93 $\pm$ 0.01 12.12 $\pm$ 0.00 50.09 $\pm$ 0.06 00.22 $\pm$ 0.00 $0$$\pm$ 0.00
VL3-2B 13.92 $\pm$ 0.13 03.47 $\pm$ 0.02 00.32 $\pm$ 0.01 13.92 $\pm$ 0.13 03.46 $\pm$ 0.04 00.32 $\pm$ 0.01
VL3-7B 15.05 $\pm$ 0.21 16.67 $\pm$ 0.35 06.35 $\pm$ 0.06 12.76 $\pm$ 0.52 05.88 $\pm$ 0.01 00.54 $\pm$ 0.01
Table 5: Performance of fine-tuned PaliGemma 2 variants on Action and Character Recognition tasks. We compare full, partial, and parameter-efficient tuning strategies. “MC” denotes multiple-choice and “OE” open-ended formats.
Binary Multiple-choice Open-ended
Model EM ROUGE EM ROUGE EM ROUGE
Action Recognition
$\mathcal{F}_{L}$ 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00
$\mathcal{F}_{P}$ 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
$\mathcal{F}_{V}$ 83.70 $\pm$ 0.97 83.70 $\pm$ 0.97 63.40 $\pm$ 0.45 69.87 $\pm$ 0.44 66.03 $\pm$ 0.10 71.92 $\pm$ 0.08
$\mathcal{F}_{P+L}$ 74.47 $\pm$ 1.64 74.47 $\pm$ 1.64 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00 55.74 $\pm$ 0.70 64.83 $\pm$ 0.29
$\mathcal{F}_{V+L}$ 75.80 $\pm$ 0.16 75.80 $\pm$ 0.16 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00
$\mathcal{F}_{V+P}$ 73.46 $\pm$ 0.85 73.46 $\pm$ 0.85 61.21 $\pm$ 0.23 67.57 $\pm$ 0.21 64.70 $\pm$ 0.02 70.93 $\pm$ 0.01
$\mathcal{F}_{all}$ 74.35 $\pm$ 1.37 74.35 $\pm$ 1.37 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00
$\mathcal{F}_{LoRA}$ 44.66 $\pm$ 0.21 44.66 $\pm$ 0.21 00.02 $\pm$ 0.01 09.21 $\pm$ 0.01 $0$$\pm$ 0.00 12.49 $\pm$ 0.00
Character Recognition
$\mathcal{F}_{L}$ 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 98.92 $\pm$ 0.00 98.92 $\pm$ 0.00 98.98 $\pm$ 0.00 98.99 $\pm$ 0.01
$\mathcal{F}_{P}$ 99.09 $\pm$ 0.11 99.09 $\pm$ 0.11 99.22 $\pm$ 0.33 99.22 $\pm$ 0.33 99.15 $\pm$ 0.07 99.15 $\pm$ 0.07
$\mathcal{F}_{V}$ 99.31 $\pm$ 0.01 99.31 $\pm$ 0.01 99.14 $\pm$ 0.42 99.14 $\pm$ 0.42 99.61 $\pm$ 0.12 99.61 $\pm$ 0.12
$\mathcal{F}_{P+L}$ 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 98.28 $\pm$ 0.00 98.30 $\pm$ 0.02 98.39 $\pm$ 0.00 98.39 $\pm$ 0.00
$\mathcal{F}_{V+L}$ 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 96.88 $\pm$ 0.00 96.88 $\pm$ 0.00 98.45 $\pm$ 0.00 98.45 $\pm$ 0.00
$\mathcal{F}_{V+P}$ 60.32 $\pm$ 0.02 60.32 $\pm$ 0.02 99.22 $\pm$ 0.00 99.22 $\pm$ 0.00 99.79 $\pm$ 0.00 99.79 $\pm$ 0.00
$\mathcal{F}_{all}$ 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 97.67 $\pm$ 0.06 97.67 $\pm$ 0.06 96.55 $\pm$ 0.01 96.55 $\pm$ 0.01
$\mathcal{F}_{LoRA}$ 48.76 $\pm$ 0.00 48.76 $\pm$ 0.00 $0$$\pm$ 0.00 00.32 $\pm$ 0.00 $0$$\pm$ 0.00 00.01 $\pm$ 0.01
Table 6: Supervision and Temporal Context: Training budget analysis for Character Recognition. Models are fine-tuned for sub-epoch durations and evaluated across binary, multiple-choice (MC), and open-ended (OE) formats.
Binary Multiple-choice Open-ended
Ep EM ROUGE EM ROUGE EM ROUGE
0.005 50.84 $\pm$ 1.70 50.84 $\pm$ 1.70 14.27 $\pm$ 0.00 14.27 $\pm$ 0.00 13.16 $\pm$ 0.00 13.27 $\pm$ 0.00
0.01 54.92 $\pm$ 0.84 54.92 $\pm$ 0.84 17.85 $\pm$ 0.00 17.85 $\pm$ 0.00 16.78 $\pm$ 0.14 16.88 $\pm$ 0.01
0.02 57.91 $\pm$ 3.19 57.91 $\pm$ 3.19 28.64 $\pm$ 0.00 28.64 $\pm$ 0.00 26.29 $\pm$ 2.96 28.38 $\pm$ 0.01
0.03 57.45 $\pm$ 0.58 57.45 $\pm$ 0.58 64.19 $\pm$ 0.00 64.19 $\pm$ 0.00 40.76 $\pm$ 0.01 40.98 $\pm$ 0.00
0.06 65.88 $\pm$ 3.71 65.88 $\pm$ 3.71 93.96 $\pm$ 0.00 93.96 $\pm$ 0.00 88.89 $\pm$ 0.00 88.95 $\pm$ 0.01
0.10 87.47 $\pm$ 4.11 87.47 $\pm$ 4.11 96.54 $\pm$ 0.00 96.54 $\pm$ 0.00 97.01 $\pm$ 0.38 97.02 $\pm$ 0.40
0.125 91.63 $\pm$ 7.71 91.63 $\pm$ 7.71 97.08 $\pm$ 0.00 97.08 $\pm$ 0.00 97.28 $\pm$ 0.00 97.30 $\pm$ 0.00
0.20 97.96 $\pm$ 0.45 97.96 $\pm$ 0.45 97.89 $\pm$ 0.28 97.89 $\pm$ 0.28 98.12 $\pm$ 0.00 98.14 $\pm$ 0.02
0.25 97.75 $\pm$ 0.74 97.75 $\pm$ 0.74 98.08 $\pm$ 0.00 98.08 $\pm$ 0.00 98.12 $\pm$ 0.00 98.15 $\pm$ 0.00
0.33 98.42 $\pm$ 0.20 98.42 $\pm$ 0.20 98.19 $\pm$ 0.00 98.19 $\pm$ 0.00 98.30 $\pm$ 0.00 98.35 $\pm$ 0.00
0.50 98.74 $\pm$ 0.06 98.74 $\pm$ 0.06 98.45 $\pm$ 0.00 98.45 $\pm$ 0.00 98.51 $\pm$ 0.00 98.54 $\pm$ 0.00
0.67 99.11 $\pm$ 0.10 99.11 $\pm$ 0.10 98.99 $\pm$ 0.08 98.99 $\pm$ 0.08 99.09 $\pm$ 0.00 99.09 $\pm$ 0.00
0.75 99.03 $\pm$ 0.04 99.03 $\pm$ 0.04 99.15 $\pm$ 0.00 99.15 $\pm$ 0.00 99.21 $\pm$ 0.00 99.22 $\pm$ 0.01
1 99.09 $\pm$ 0.11 99.09 $\pm$ 0.11 99.22 $\pm$ 0.33 99.22 $\pm$ 0.33 99.15 $\pm$ 0.07 99.15 $\pm$ 0.07
Table 7: Supervision and Temporal Context: Training budget analysis for Action Recognition, with models fine-tuned for up to 10 epochs. Evaluated using across binary, multiple-choice (MC), and open-ended (OE) formats.
Binary Multiple-Choice Open-Ended
Ep EM ROUGE EM ROUGE EM ROUGE
1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
2 84.92 $\pm$ 0.23 84.92 $\pm$ 0.23 64.90 $\pm$ 0.15 71.17 $\pm$ 0.01 64.05 $\pm$ 0.01 70.36 $\pm$ 0.01
4 85.37 $\pm$ 0.30 85.37 $\pm$ 0.30 64.58 $\pm$ 0.69 70.89 $\pm$ 0.55 64.79 $\pm$ 0.31 70.88 $\pm$ 0.27
8 85.18 $\pm$ 0.20 85.18 $\pm$ 0.20 63.53 $\pm$ 0.35 69.95 $\pm$ 0.37 63.43 $\pm$ 0.28 69.91 $\pm$ 0.18
10 85.11 $\pm$ 0.41 85.11 $\pm$ 0.41 62.88 $\pm$ 1.35 69.34 $\pm$ 1.20 66.82 $\pm$ 3.75 72.34 $\pm$ 3.04
Table 8: Supervision and Temporal Context: Training budget and temporal context analysis for Action Recognition. Models are fine-tuned for up to 10 epochs and evaluated with up to 8 input frames.
Binary Multiple-choice Open-ended
Ep Fr EM ROUGE EM ROUGE EM ROUGE
1 1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
2 84.42 $\pm$ 0.06 84.42 $\pm$ 0.06 65.53 $\pm$ 0.27 72.03 $\pm$ 0.06 65.38 $\pm$ 0.06 71.74 $\pm$ 0.23
4 90.97 $\pm$ 0.10 90.97 $\pm$ 0.10 83.11 $\pm$ 0.08 87.13 $\pm$ 0.04 82.26 $\pm$ 0.14 87.02 $\pm$ 0.06
8 93.85 $\pm$ 0.28 93.85 $\pm$ 0.28 88.89 $\pm$ 0.14 93.40 $\pm$ 0.76 87.80 $\pm$ 0.20 92.23 $\pm$ 0.16
2 1 85.10 $\pm$ 0.02 85.10 $\pm$ 0.02 64.93 $\pm$ 0.06 71.09 $\pm$ 0.11 64.05 $\pm$ 0.01 70.36 $\pm$ 0.01
2 86.53 $\pm$ 0.45 86.40 $\pm$ 0.26 69.20 $\pm$ 0.88 75.02 $\pm$ 0.67 68.83 $\pm$ 0.47 72.45 $\pm$ 2.76
4 92.26 $\pm$ 0.34 92.26 $\pm$ 0.34 84.19 $\pm$ 0.10 88.46 $\pm$ 0.11 83.34 $\pm$ 0.15 87.84 $\pm$ 0.06
8 95.05 $\pm$ 0.15 95.05 $\pm$ 0.15 89.27 $\pm$ 0.14 93.30 $\pm$ 0.22 89.42 $\pm$ 0.22 93.25 $\pm$ 0.15
4 1 85.37 $\pm$ 0.30 85.37 $\pm$ 0.30 64.58 $\pm$ 0.69 70.89 $\pm$ 0.55 64.79 $\pm$ 0.31 70.88 $\pm$ 0.27
2 86.89 $\pm$ 0.06 86.89 $\pm$ 0.06 69.89 $\pm$ 0.83 75.49 $\pm$ 0.69 70.04 $\pm$ 0.08 75.74 $\pm$ 0.06
4 92.58 $\pm$ 0.18 92.58 $\pm$ 0.18 85.13 $\pm$ 0.19 89.28 $\pm$ 0.17 84.61 $\pm$ 0.07 88.81 $\pm$ 0.04
8 95.29 $\pm$ 0.07 95.29 $\pm$ 0.07 90.64 $\pm$ 0.00 94.09 $\pm$ 0.04 90.18 $\pm$ 0.16 93.81 $\pm$ 0.11
8 1 85.04 $\pm$ 0.00 85.04 $\pm$ 0.00 63.53 $\pm$ 0.35 69.95 $\pm$ 0.37 63.62 $\pm$ 0.00 70.03 $\pm$ 0.00
2 87.27 $\pm$ 0.40 87.27 $\pm$ 0.40 69.84 $\pm$ 0.66 75.44 $\pm$ 0.45 70.11 $\pm$ 0.65 75.75 $\pm$ 0.52
4 92.97 $\pm$ 0.49 92.97 $\pm$ 0.49 85.32 $\pm$ 0.08 89.27 $\pm$ 0.08 84.93 $\pm$ 0.21 89.05 $\pm$ 0.11
8 95.48 $\pm$ 0.21 95.48 $\pm$ 0.21 90.71 $\pm$ 0.14 94.15 $\pm$ 0.13 91.02 $\pm$ 0.28 93.96 $\pm$ 0.71
10 1 85.40 $\pm$ 0.00 85.40 $\pm$ 0.00 62.88 $\pm$ 1.35 69.34 $\pm$ 1.20 66.82 $\pm$ 3.75 72.34 $\pm$ 3.04
2 87.17 $\pm$ 0.22 87.17 $\pm$ 0.22 70.18 $\pm$ 0.00 75.64 $\pm$ 0.00 69.59 $\pm$ 0.20 75.37 $\pm$ 0.07
4 92.96 $\pm$ 0.37 92.96 $\pm$ 0.37 85.02 $\pm$ 0.58 89.05 $\pm$ 0.49 84.71 $\pm$ 0.08 88.94 $\pm$ 0.06
8 96.03 $\pm$ 0.05 96.03 $\pm$ 0.05 90.75 $\pm$ 0.04 94.22 $\pm$ 0.05 91.00 $\pm$ 0.11 94.33 $\pm$ 0.09
Table 9: Comparison of frame sampling strategies for Action Recognition. We evaluate first-$n$ vs. uniform-$n$ sampling across varying temporal context lengths ($n\in\{1,\dots,8\}$).
Binary Multiple-choice Open-ended
Fr EM ROUGE EM ROUGE EM ROUGE
First-N 1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
2 84.42 $\pm$ 0.06 84.42 $\pm$ 0.06 65.53 $\pm$ 0.27 72.03 $\pm$ 0.06 65.38 $\pm$ 0.06 71.74 $\pm$ 0.23
3 87.93 $\pm$ 0.28 87.93 $\pm$ 0.28 75.73 $\pm$ 0.18 81.07 $\pm$ 0.06 74.68 $\pm$ 0.16 80.21 $\pm$ 0.08
4 90.97 $\pm$ 0.10 90.97 $\pm$ 0.10 83.11 $\pm$ 0.08 87.13 $\pm$ 0.04 82.26 $\pm$ 0.14 87.02 $\pm$ 0.06
5 92.00 $\pm$ 0.30 92.00 $\pm$ 0.30 85.46 $\pm$ 0.34 89.84 $\pm$ 0.18 85.10 $\pm$ 0.16 89.47 $\pm$ 0.10
6 92.95 $\pm$ 0.30 92.95 $\pm$ 0.30 86.86 $\pm$ 0.08 91.13 $\pm$ 0.06 86.59 $\pm$ 0.39 90.82 $\pm$ 0.30
7 93.31 $\pm$ 0.03 93.31 $\pm$ 0.03 87.95 $\pm$ 0.08 92.06 $\pm$ 0.06 87.58 $\pm$ 0.17 91.82 $\pm$ 0.08
8 93.85 $\pm$ 0.28 93.85 $\pm$ 0.28 88.89 $\pm$ 0.14 93.40 $\pm$ 0.76 87.80 $\pm$ 0.20 92.23 $\pm$ 0.16
Uniform-N 1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
2 90.47 $\pm$ 0.62 90.47 $\pm$ 0.62 83.93 $\pm$ 0.04 88.36 $\pm$ 0.08 82.68 $\pm$ 0.19 87.33 $\pm$ 0.01
3 93.59 $\pm$ 0.07 93.59 $\pm$ 0.07 88.90 $\pm$ 0.11 92.85 $\pm$ 0.10 88.49 $\pm$ 0.24 92.57 $\pm$ 0.04
4 93.57 $\pm$ 0.39 93.57 $\pm$ 0.39 89.94 $\pm$ 0.04 93.65 $\pm$ 0.01 89.56 $\pm$ 0.42 93.49 $\pm$ 0.28
5 94.25 $\pm$ 0.04 94.25 $\pm$ 0.04 89.99 $\pm$ 0.18 93.70 $\pm$ 0.13 89.72 $\pm$ 0.10 93.63 $\pm$ 0.23
6 94.01 $\pm$ 0.57 94.01 $\pm$ 0.57 90.03 $\pm$ 0.04 93.73 $\pm$ 0.06 90.09 $\pm$ 0.18 93.88 $\pm$ 0.16
7 93.96 $\pm$ 0.16 93.96 $\pm$ 0.16 90.34 $\pm$ 0.23 94.00 $\pm$ 0.10 89.94 $\pm$ 0.11 93.73 $\pm$ 0.10
8 94.62 $\pm$ 0.48 94.62 $\pm$ 0.48 90.72 $\pm$ 0.12 94.30 $\pm$ 0.10 90.30 $\pm$ 0.04 94.01 $\pm$ 0.02
Table 10: Optimizing Data Mix for Unified Multi-Task Evaluation: Performance tradeoffs under varying task-level allocation ratios for Action ($\alpha_{\text{AR}}$) vs. Character Recognition ($\alpha_{\text{CR}}$), with a fixed format distribution ($\beta=\nicefrac{{1}}{{3}}$ per format). Evaluated across binary, multiple-choice (MC), and open-ended (OE) formats.
Binary Multiple-choice Open-ended
$\alpha_{\text{AR}}$ $\alpha_{\text{CR}}$ EM ROUGE EM ROUGE EM ROUGE
Action Recognition
0.20 0.80 84.13 $\pm$ 1.66 84.13 $\pm$ 1.66 26.10 $\pm$ 0.43 39.04 $\pm$ 0.10 27.03 $\pm$ 0.21 39.60 $\pm$ 0.13
0.40 0.60 88.11 $\pm$ 1.44 88.11 $\pm$ 1.44 28.13 $\pm$ 1.15 40.93 $\pm$ 0.57 29.17 $\pm$ 0.40 41.66 $\pm$ 0.51
[…]
0.60 0.40 90.80 $\pm$ 0.04 90.80 $\pm$ 0.04 30.44 $\pm$ 0.03 42.54 $\pm$ 0.01 30.55 $\pm$ 0.49 42.32 $\pm$ 0.60
0.80 0.20 91.23 $\pm$ 0.91 91.23 $\pm$ 0.91 84.06 $\pm$ 1.26 89.42 $\pm$ 1.00 30.88 $\pm$ 0.63 42.85 $\pm$ 0.42
Character Recognition
0.20 0.80 98.57 $\pm$ 0.47 98.57 $\pm$ 0.47 98.95 $\pm$ 0.16 98.95 $\pm$ 0.16 98.94 $\pm$ 0.22 98.97 $\pm$ 0.21
0.40 0.60 98.51 $\pm$ 0.53 98.51 $\pm$ 0.53 98.77 $\pm$ 0.16 98.77 $\pm$ 0.16 98.98 $\pm$ 0.06 98.98 $\pm$ 0.06
[…]
0.60 0.40 93.22 $\pm$ 3.38 93.22 $\pm$ 3.38 96.91 $\pm$ 1.81 96.94 $\pm$ 1.77 97.93 $\pm$ 0.02 97.93 $\pm$ 0.02
0.80 0.20 80.53 $\pm$ 0.49 80.53 $\pm$ 0.49 89.08 $\pm$ 0.49 89.08 $\pm$ 0.49 89.02 $\pm$ 0.25 89.18 $\pm$ 0.39
Table 11: Optimizing Data Mix for Unified Multi-Task Evaluation: Performance on Action and Character Recognition under varying format-level sampling ratios ($\beta$) for Binary, Multiple-choice (MC), and Open-ended (OE) questions. We fix $\alpha_{\texttt{AR}}=0.8$ and train all models on the first 8 frames.
Binary Multiple-Choice Open-Ended
Ep $\beta_{\texttt{binary}}$ $\beta_{\texttt{MC}}$ $\beta_{\texttt{OE}}$ EM ROUGE EM ROUGE EM ROUGE
Action Recognition
1 0.4 0.2 0.4 92.32 $\pm$ 0.37 92.32 $\pm$ 0.37 84.32 $\pm$ 0.89 89.54 $\pm$ 0.64 31.61 $\pm$ 0.23 43.51 $\pm$ 0.09
0.2 0.4 0.4 90.80 $\pm$ 0.71 90.80 $\pm$ 0.71 86.60 $\pm$ 0.38 91.27 $\pm$ 0.42 32.06 $\pm$ 0.18 43.58 $\pm$ 0.54
0.0 0.4 0.6 49.98 $\pm$ 0.04 49.98 $\pm$ 0.04 86.65 $\pm$ 0.99 91.26 $\pm$ 0.87 85.51 $\pm$ 1.78 90.13 $\pm$ 1.53
0.0 0.2 0.8 50.11 $\pm$ 0.06 50.11 $\pm$ 0.06 86.58 $\pm$ 0.47 91.42 $\pm$ 0.21 87.45 $\pm$ 0.09 91.78 $\pm$ 0.09
2 0.4 0.2 0.4 93.14 $\pm$ 0.48 93.14 $\pm$ 0.48 86.77 $\pm$ 0.00 91.28 $\pm$ 0.00 32.96 $\pm$ 0.00 44.00 $\pm$ 0.00
0.2 0.4 0.4 92.89 $\pm$ 0.16 92.89 $\pm$ 0.16 87.83 $\pm$ 0.00 92.13 $\pm$ 0.00 33.37 $\pm$ 0.00 44.13 $\pm$ 0.00
0.0 0.4 0.6 41.22 $\pm$ 0.00 41.30 $\pm$ 0.01 89.17 $\pm$ 0.07 93.12 $\pm$ 0.02 88.55 $\pm$ 0.00 93.65 $\pm$ 0.00
0.0 0.2 0.8 49.98 $\pm$ 0.03 49.99 $\pm$ 0.02 88.68 $\pm$ 0.00 92.71 $\pm$ 0.00 88.59 $\pm$ 0.00 92.56 $\pm$ 0.00
4 0.4 0.2 0.4 94.33 $\pm$ 0.34 94.33 $\pm$ 0.34 92.67 $\pm$ 0.07 93.27 $\pm$ 0.71 33.94 $\pm$ 0.01 43.96 $\pm$ 0.00
0.2 0.4 0.4 94.19 $\pm$ 0.13 94.19 $\pm$ 0.13 93.04 $\pm$ 0.01 93.57 $\pm$ 0.74 33.78 $\pm$ 0.00 44.73 $\pm$ 0.00
0.0 0.4 0.6 50.19 $\pm$ 0.27 50.19 $\pm$ 0.27 89.78 $\pm$ 0.11 93.52 $\pm$ 0.03 88.95 $\pm$ 0.05 92.75 $\pm$ 0.06
0.0 0.2 0.8 49.98 $\pm$ 0.02 49.98 $\pm$ 0.02 89.25 $\pm$ 0.00 93.13 $\pm$ 0.00 89.41 $\pm$ 0.00 93.16 $\pm$ 0.00
Character Recognition
1 0.4 0.2 0.4 86.42 $\pm$ 0.25 86.42 $\pm$ 0.25 94.77 $\pm$ 0.14 94.77 $\pm$ 0.14 94.76 $\pm$ 0.08 94.63 $\pm$ 0.21
0.2 0.4 0.4 77.57 $\pm$ 0.01 77.57 $\pm$ 0.01 94.93 $\pm$ 0.03 94.93 $\pm$ 0.03 94.51 $\pm$ 0.57 94.15 $\pm$ 0.00
0.0 0.4 0.6 50.37 $\pm$ 0.52 50.37 $\pm$ 0.52 96.56 $\pm$ 0.28 96.56 $\pm$ 0.28 96.81 $\pm$ 0.39 96.82 $\pm$ 0.37
0.0 0.2 0.8 50.51 $\pm$ 0.03 50.51 $\pm$ 0.03 96.88 $\pm$ 0.27 96.88 $\pm$ 0.27 97.39 $\pm$ 0.18 97.39 $\pm$ 0.18
2 0.4 0.2 0.4 89.95 $\pm$ 0.46 89.95 $\pm$ 0.46 87.35 $\pm$ 0.00 87.36 $\pm$ 0.00 88.64 $\pm$ 0.00 88.64 $\pm$ 0.00
0.2 0.4 0.4 91.28 $\pm$ 0.20 91.28 $\pm$ 0.20 93.90 $\pm$ 0.00 93.93 $\pm$ 0.04 93.06 $\pm$ 0.00 93.09 $\pm$ 0.00
0.0 0.4 0.6 47.37 $\pm$ 0.02 47.48 $\pm$ 0.04 97.69 $\pm$ 0.00 97.70 $\pm$ 0.00 98.07 $\pm$ 0.00 98.07 $\pm$ 0.00
0.0 0.2 0.8 50.07 $\pm$ 0.04 50.07 $\pm$ 0.04 97.75 $\pm$ 0.00 97.79 $\pm$ 0.00 98.07 $\pm$ 0.00 98.07 $\pm$ 0.00
4 0.4 0.2 0.4 97.71 $\pm$ 0.05 97.71 $\pm$ 0.05 97.70 $\pm$ 0.00 97.70 $\pm$ 0.00 97.88 $\pm$ 0.01 97.91 $\pm$ 0.03
0.2 0.4 0.4 96.55 $\pm$ 0.06 96.55 $\pm$ 0.06 98.51 $\pm$ 0.00 98.51 $\pm$ 0.00 98.46 $\pm$ 0.00 98.46 $\pm$ 0.00
0.0 0.4 0.6 51.25 $\pm$ 1.56 51.25 $\pm$ 1.56 98.21 $\pm$ 0.13 98.21 $\pm$ 0.13 98.48 $\pm$ 0.06 98.49 $\pm$ 0.06
0.0 0.2 0.8 50.93 $\pm$ 0.10 50.93 $\pm$ 0.10 98.79 $\pm$ 0.00 98.79 $\pm$ 0.00 99.08 $\pm$ 0.00 99.09 $\pm$ 0.00
Table 12: Optimizing Data Mix for Unified Multi-Task Evaluation: Performance on Action and Character Recognition under high open-ended (OE) supervision, with $\beta_{\texttt{OE}}=0.8$ and remaining budget split between Binary and Multiple-choice (MC).
Binary Multiple-Choice Open-Ended
Ep $\beta_{\texttt{BN}}$ $\beta_{\texttt{MC}}$ $\beta_{\texttt{OE}}$ EM ROUGE EM ROUGE EM ROUGE
Action Recognition
1 0.15 0.05 0.80 88.85 $\pm$ 0.04 88.85 $\pm$ 0.04 81.93 $\pm$ 0.00 89.08 $\pm$ 0.00 86.72 $\pm$ 0.00 91.33 $\pm$ 0.00
0.10 0.10 0.80 87.38 $\pm$ 0.59 87.38 $\pm$ 0.59 85.58 $\pm$ 0.00 90.50 $\pm$ 0.00 86.88 $\pm$ 0.00 91.25 $\pm$ 0.00
0.05 0.15 0.80 85.67 $\pm$ 0.19 85.67 $\pm$ 0.19 86.38 $\pm$ 0.09 91.21 $\pm$ 0.06 86.84 $\pm$ 0.02 91.34 $\pm$ 0.03
2 0.15 0.05 0.80 92.45 $\pm$ 0.11 92.45 $\pm$ 0.11 87.52 $\pm$ 0.00 91.90 $\pm$ 0.00 87.97 $\pm$ 0.63 92.19 $\pm$ 0.41
0.10 0.10 0.80 92.11 $\pm$ 0.18 92.11 $\pm$ 0.18 88.42 $\pm$ 0.00 92.54 $\pm$ 0.00 88.50 $\pm$ 0.00 92.54 $\pm$ 0.00
0.05 0.15 0.80 91.98 $\pm$ 0.24 91.98 $\pm$ 0.24 88.72 $\pm$ 0.00 92.78 $\pm$ 0.00 88.56 $\pm$ 0.00 92.66 $\pm$ 0.00
4 0.15 0.05 0.80 92.98 $\pm$ 0.21 92.98 $\pm$ 0.21 88.93 $\pm$ 0.00 93.02 $\pm$ 0.00 89.64 $\pm$ 0.00 93.34 $\pm$ 0.00
0.10 0.10 0.80 92.81 $\pm$ 0.11 92.81 $\pm$ 0.11 91.40 $\pm$ 2.81 93.88 $\pm$ 0.70 89.43 $\pm$ 0.00 93.20 $\pm$ 0.00
0.05 0.15 0.80 91.52 $\pm$ 0.37 91.52 $\pm$ 0.37 89.80 $\pm$ 0.00 93.54 $\pm$ 0.01 89.81 $\pm$ 0.01 93.49 $\pm$ 0.06
Character Recognition
1 0.15 0.05 0.80 59.75 $\pm$ 0.04 59.75 $\pm$ 0.04 95.45 $\pm$ 0.00 95.45 $\pm$ 0.00 97.16 $\pm$ 0.00 97.16 $\pm$ 0.00
0.10 0.10 0.80 56.31 $\pm$ 0.53 56.31 $\pm$ 0.53 94.55 $\pm$ 0.00 94.55 $\pm$ 0.00 95.96 $\pm$ 0.00 95.96 $\pm$ 0.00
0.05 0.15 0.80 50.86 $\pm$ 0.08 50.86 $\pm$ 0.08 96.87 $\pm$ 0.02 96.87 $\pm$ 0.02 97.12 $\pm$ 0.01 97.12 $\pm$ 0.01
2 0.15 0.05 0.80 80.18 $\pm$ 0.09 80.18 $\pm$ 0.09 95.41 $\pm$ 0.00 95.41 $\pm$ 0.00 96.91 $\pm$ 0.00 96.91 $\pm$ 0.00
0.10 0.10 0.80 70.20 $\pm$ 0.21 70.20 $\pm$ 0.21 98.02 $\pm$ 0.00 98.02 $\pm$ 0.00 98.15 $\pm$ 0.00 98.15 $\pm$ 0.00
0.05 0.15 0.80 69.67 $\pm$ 0.36 69.67 $\pm$ 0.36 97.37 $\pm$ 0.00 97.37 $\pm$ 0.00 97.79 $\pm$ 0.00 97.79 $\pm$ 0.00
4 0.15 0.05 0.80 94.16 $\pm$ 0.12 94.16 $\pm$ 0.12 97.06 $\pm$ 0.00 97.07 $\pm$ 0.00 97.91 $\pm$ 0.00 97.91 $\pm$ 0.00
0.10 0.10 0.80 86.67 $\pm$ 0.13 86.67 $\pm$ 0.13 98.50 $\pm$ 0.00 98.50 $\pm$ 0.00 98.77 $\pm$ 0.00 98.77 $\pm$ 0.00
0.05 0.15 0.80 71.79 $\pm$ 0.01 71.79 $\pm$ 0.01 98.22 $\pm$ 0.00 98.22 $\pm$ 0.00 98.57 $\pm$ 0.00 98.57 $\pm$ 0.00
Appendix E Human Annotation Study
This section provides full details of our human annotation study, including rollout generation, annotation procedures, inter-annotator agreement, and evaluation metrics. The goal is to validate the adapted VLM’s fine-grained predictions on generated video rollouts.
E.1 Study Design
Table 13: Annotation instructions provided to human raters as part of the study. The interface outlines task context, scoring criteria, general guidelines, and reference definitions for supported action categories.
[…]
You will be presented with: A short video clip; A natural language question about the video; An answer generated by a vision-language model. Your task is to evaluate whether the model’s answer accurately describes the events depicted in the video.
[…]
Assign one of the following categories: Correct (1.0): Fully matches the event in the video; Partially Correct (0.5): Captures the general idea but contains a minor error; Incorrect (0.0): Wrong, hallucinated, or mismatched with the visual evidence; Unclear / Cannot Tell: Not enough evidence to confidently decide.
[…]
Watch the full video before rating; Base your decision solely on visible content; Use provided action and character references; If multiple plausible interpretations exist and the answer matches one, mark as Correct; If unsure even after review, mark Unclear / Cannot Tell; Optionally leave comments for ambiguous or interesting cases.
[…]
Evading Backwards: Moves backwards to avoid threat or reposition. Evading Forwards: Moves forwards. Evading Left / Right: Lateral movement left or right. Jumping Down: Jumps from a higher to a lower platform or level. Jumping on the Level: Jumps without elevation change. Jumping Up: Jumps upward to reach a higher platform. Mounting Hoverboard: Begins riding or is seen riding a hoverboard.
Figure 9: Reference slides shown to annotators during the human annotation study, illustrating the two recognition targets: actions (left) and characters (center and right). The slides include 20 exemplar videos (7 actions, 13 characters) to support consistent evaluation of VLM-generated responses.
Task Overview. Human annotators were presented with short video clips generated by a world model, each paired with a natural language question and an answer generated by the VLM. They were asked to judge whether the model’s answer accurately described what was shown in the video. Each QA pair was rated using one of four categories: Correct (score = 1), Partially Correct (0.5), Incorrect (0), or Unclear / Cannot Tell (excluded from accuracy computation).
Annotation Setup and Interface. Annotations were collected using a custom PowerPoint-based interface (see Figure 10). Each slide presented a short video, a question, and a generated answer. Annotators selected a rating from a predefined rubric. The full annotation guidelines – including action and character definitions and rating instructions – were embedded in the annotation deck for reference. For completeness, we also provide them in Table 13 and Figure 9. The annotation study was carried out by a subset of the authors with prior experience in the environment. Judging correctness required non-trivial familiarity with the visual dynamics and task ontology, making expert annotation necessary. All annotators were compensated above local minimum wage rates. Each QA pair was independently rated by two primary annotators. In cases of disagreement or if either annotator marked the example as Unclear, a third, more experienced adjudicator reviewed the pair and assigned a final rating.
Selected World Models. For our study, we select two autoregressive world models of different scales, both based on the WHAM architecture [56], a publicly available world model. These models are trained to model sequences of visual frames and controller actions, without any textual supervision. Each world model is a decoder-only transformer [81, 98] trained to autoregressively predict discrete tokens representing visual observations and actions. Visual frames are first encoded using a VQGAN [32], while joystick actions are tokenized using a learned discretization scheme based on action bucketization [55]. The model is trained to predict the next token in the sequence, conditioned on prior visual and action tokens. Specifically, we focus on two versions of WHAM with differing model capacities and training environments:
  • WHAM 140M: A 140M-parameter model trained for 100K steps on gameplay from a single environment (Environment A / Skygarden) at $128\times 128$ resolution.
  • WHAM 1.6B: A 1.6B-parameter model trained for 200K steps on gameplay from seven environments (Environments A–G, including Skygarden) at $300\times 180$ resolution.
Rollouts Generation. Rollout generation follows a consistent protocol for both world models: at inference time, the model is conditioned on 1 second of ground-truth gameplay (visual and action tokens), after which it generates 10 seconds of future gameplay conditioned only on a sequence of held-out controller actions. The generated rollout is then split into 14-frame chunks. This setup enables a comprehensive analysis of the universe’s evaluation capabilities across two axes: (i) in-domain performance: evaluating on Skygarden (Environment A), the environment used for fine-tuning; (ii) generalization: assessing performance on six unseen environments (Environments B–G). It also allows comparison across generation quality and model capacity. We generate 82 rollouts for each model-environment setting, resulting in 656 rollouts in total.
Figure 10: Annotation interface example. Each instance includes a video clip, task instructions, and a table with: Question (generated via evaluation protocol), Response (VLM output), and Score (human-assigned label).
Rollout Filtering. To ensure quality and clarity, we filtered out rollouts that: (i) had no visible agents, (ii) featured stationary agents, (iii) were taken from early uninformative environment segments, or (iv) had significant visual obstruction. We also excluded sequences containing more than four characters to reduce annotation ambiguity.
universe Response Generation. To obtain responses from universe, we provide it with a video segment (resized to match the evaluator’s input resolution) along with its corresponding question. We then sample five responses using greedy decoding and select the most frequent response as the final answer. In cases where all five responses are unique (i.e., no majority), one response is selected at random. The resulting dataset comprises rollouts from 8 model–environment pairs: rollouts generated by WHAM 140M on Environment A (Skygarden), and rollouts generated by WHAM 1.6B across seven distinct environments (Environments A–G). For each model–environment pair, we sample 30 rollouts. Each rollout is annotated with 6 question–answer (QA) pairs, along with a corresponding response from the adapted evaluator. Each of the resulting 1,440 QA instances was rated by 3 annotators, yielding 4,320 total human judgments.
E.2 Evaluation Metrics
We report two accuracy-based metrics using the adjudicated labels:
Strict Accuracy.: The proportion of QA pairs labeled as Correct:
$\text{Acc}_{\text{Strict}}=\frac{N_{\text{Correct}}}{N_{\text{Answerable}}},$ (5)
Graded Accuracy.: Partial credit given to Partially Correct responses:
$\text{Acc}_{\text{Graded}}=\frac{N_{\text{Correct}}+0.5\times N_{\text{Partial}}}{N_{\text{Answerable}}}.$ (6)
Only examples not marked Unclear by adjudication are included in $N_{\text{Answerable}}$.
Inter-Annotator Agreement. To quantify rating consistency, we compute Cohen’s $\kappa$ between the two primary annotators. The adjudicator’s label is used only when disagreement occurs and is excluded from agreement computation. Results are shown in Table 14.
Table 14: Inter-annotator agreement and valid QA coverage across environments. We report Cohen’s $\kappa$ between the two primary annotators for each world model–map pair. The total number of valid examples excludes QA pairs marked as Unclear by at least one annotator.
Model Env. Valid QA Pairs Cohen’s $\kappa$
WHAM 140M A 24 0.79
WHAM 1.6B A 29 0.91
B 28 0.67
C 28 0.74
D 29 0.87
E 30 0.67
F 29 0.61
G 30 0.59
Sample Size Justification. We annotate 30 rollouts per model–environment pair. Assuming a standard deviation of $\sigma\approx 0.2$ and a 95% confidence level, the confidence interval (CI) width is given by CI $\text{CI Width}=z_{\frac{1-C}{2}}\cdot\frac{\sigma}{\sqrt{n}}$. This yields an estimated CI of $\sim$7.1for individual model–environment pairs ($n=30$), and $\sim$2.5when aggregating across all eight pairs ($n=240$), offering sufficient precision for comparative evaluation.
Table 15: Graded/strict accuracy of universe on Action and Character Recognition tasks, evaluated by human annotators across different environments and question formats. We report results for Binary, Multiple-Choice (MC), and Open-Ended (OE) prompts, disaggregated by task and world model. All metrics are based on final adjudicated ratings.
Action Recognition Character Recognition
Model Env. Binary MC OE Binary MC OE
WHAM 140M A 92.9 / 89.3 35.7 / 32.1 41.1 / 39.3 85.7 / 85.7 10.7 / 10.7 60.7 / 60.7
WHAM 1.6B A 98.3 / 96.7 51.7 / 46.7 75.0 / 73.3 93.3 / 93.3 83.3 / 83.3 93.3 / 93.3
B 96.7 / 96.7 60.0 / 60.0 65.0 / 60.0 99.9 / 99.9 90.0 / 90.0 93.3 / 93.3
C 96.7 / 96.7 63.3 / 63.3 80.0 / 80.0 99.9 / 99.9 86.7 / 86.7 93.3 / 93.3
D 93.3 / 93.3 43.3 / 43.3 73.3 / 73.3 96.7 / 96.7 96.7 / 96.7 99.9 / 99.9
E 80.0 / 76.7 76.7 / 73.3 93.3 / 93.3 96.7 / 96.7 99.9 / 99.9 99.8 / 99.8
F 71.7 / 70.0 56.7 / 56.7 75.0 / 70.0 96.7 / 96.7 93.3 / 93.3 96.7 / 96.7
G 68.3 / 66.7 50.0 / 46.7 80.0 / 76.7 93.3 / 93.3 90.0 / 90.0 96.7 / 96.7
E.3 Results
[…]
Qualitative Examples. Figure 11 illustrates the diversity of generated rollouts across environments. WHAM 1.6B captures greater visual variation and scene composition compared to WHAM 140M.
Figure 11: Representative frames from rollouts generated across seven environments. WHAM 140M (top row) was trained only on Skygarden (Environment A); WHAM 1.6B (rows 2-8) generalizes across seven environments.
Appendix F Supplementary Experimental Results
This section presents additional experimental results that support the main findings but are omitted from the main paper for clarity and space. These include: (i) a zero-shot analysis of PaliGemma variants to motivate backbone selection, (ii) CLIPScore-based baselines to contextualize performance without adaptation, and (iii) a study of low-rank adaptation (LoRA) across different rank values. While these results are not central to the unified evaluation framework proposed in the main text, they provide valuable insight into model selection, adaptation efficiency, and the limitations of standard evaluation proxies in our setting.
F.1 Zero-Shot Performance of PaliGemma Models
Figure 12: Zero-shot evaluation results for PaliGemma variants across tasks, prompt formats, and visual context sizes. Overall performance remains limited, indicating the need for task-specific adaptation.
In this section, we benchmark three pretrained configurations—PaliGemma 1 3b, PaliGemma 2 3b, and PaliGemma 2 10b—under our proposed protocol and motivate our choice of PaliGemma 2 3b as the default backbone for subsequent experiments. Each model receives a natural language prompt along with either 1 or 8 image frames as input and produces a textual response. This experiment probes both model capacity and the role of temporal visual context in zero-shot settings.
Results. Figure 12 reports ROUGE scores across task types, question formats, and visual context lengths. While zero-shot performance reveals some capacity for structured reasoning—particularly in the multiple-choice setting—it remains limited overall. Binary accuracy hovers near chance, and open-ended responses frequently lack specificity. Performance is strongest on action recognition (AR), likely reflecting pretrained models’ familiarity with generic visual dynamics. In contrast, character recognition (CR) lags behind, underscoring a lack of grounding in domain-specific entities. Increasing the number of input frames modestly improves AR, but yields diminishing returns for CR. Among the evaluated configurations, PaliGemma 2 10b performs best in absolute terms. However, the margin over PaliGemma 2 3b is narrow, and PaliGemma 2 3b offers a substantially smaller footprint while using a newer Gemma 2 decoder architecture. We therefore adopt PaliGemma 2 3b as the default model for all subsequent adaptation experiments, balancing performance, compute efficiency, and architectural recency.
F.2 CLIPScore Comparisons
Table 16: Zero-shot accuracy-based evaluation of CLIP models and baseline methods on Action and Character Recognition tasks using 1 and 8 input frames.
Fr Model Action Recognition Character Recognition
1 CLIP ViT-B/32 24.04 $\pm$ 0.00 13.32 $\pm$ 0.00
CLIP ViT-B/16 52.67 $\pm$ 0.00 16.47 $\pm$ 0.00
CLIP ViT-L/14 24.60 $\pm$ 0.00 09.95 $\pm$ 0.00
CLIP ViT-L/14-336 12.17 $\pm$ 0.00 08.85 $\pm$ 0.05
8 CLIP ViT-B/32 36.22 $\pm$ 0.00 14.41 $\pm$ 0.00
CLIP ViT-B/16 57.36 $\pm$ 0.00 17.24 $\pm$ 0.00
CLIP ViT-L/14 17.57 $\pm$ 0.00 10.10 $\pm$ 0.00
CLIP ViT-L/14-336 23.12 $\pm$ 0.00 08.64 $\pm$ 0.00
To further evaluate zero-shot recognition capabilities without adaptation, we apply CLIPScore to our rollout evaluation protocol. Specifically, we assess four pretrained CLIP variants – ViT-B/32, ViT-B/16, ViT-L/14, and ViT-L/14-336 – across both Action Recognition (AR) and Character Recognition (CR) tasks using 1-frame and 8-frame visual inputs. For each evaluation instance, we extract either 1 or 8 frames from the video segment and compute the cosine similarity between each image and a predefined set of textual labels (i.e., action verbs for AR, character names for CR). For single-frame settings, we select the label with the highest similarity score as the predicted class. In the multi-frame setting, we compute predictions for each frame independently and use a majority vote to produce the final prediction. We also report two reference baselines for context: a random classifier, which achieves 12.5% on AR and 7.7% on CR, and a majority-class predictor, which yields 35.5% and 17.6% respectively. These are included only for calibration.
Results. Table 16 demonstrates the results. While CLIP ViT-B/16 performs relatively well on AR in both input settings, performance remains inconsistent across model scales and tasks. In particular, CR accuracy remains low, reflecting CLIP’s limited grounding in domain-specific visual semantics and fine-grained identity resolution. Larger CLIP models such as ViT-L/14 do not consistently outperform smaller variants, and 8-frame inputs provide only marginal gains over single-frame inputs.
Overall, these results suggest that while CLIPScore offers a lightweight and scalable evaluation proxy, it lacks the temporal grounding and semantic specificity required for structured rollout evaluation. Performance falls short relative to our selected baselines, and the method is inherently constrained to predefined candidate sets—limiting its applicability to open-ended or compositional tasks. As such, we exclude CLIP-based scores from our primary comparisons and instead focus on adapted, generative VLM-based evaluators.
F.3 Low-Rank Adaptation Comparisons
This section presents an extended analysis of low-rank adaptation (LoRA) as a parameter-efficient strategy for adapting vision-language models to our protocol. We systematically vary the rank parameter $r$ and measure its impact on Action and Character Recognition performance across all prompt formats. All experiments in this section are conducted using PaliGemma 2 (3B) as the backbone model, consistent with the main fine-tuning results. These experiments assess whether increasing rank provides meaningful gains, and inform our decision to report only the rank-8 setting in the main paper.
Table 17: Performance on Action and Character Recognition tasks after LoRA-based adaptation with varying ranks ($r\in\{8,16,32,48,64\}$). Adapters are applied to attention and MLP layers in both vision and language components.
Binary Multiple-choice Open-ended
Rank EM ROUGE EM ROUGE EM ROUGE
Action Recognition
8 44.66 $\pm$ 0.21 44.66 $\pm$ 0.21 0.02 $\pm$ 0.00 9.21 $\pm$ 0.00 0.00 $\pm$ 0.00 12.49 $\pm$ 0.00
16 44.47 $\pm$ 0.43 44.47 $\pm$ 0.43 0.02 $\pm$ 0.00 9.21 $\pm$ 0.00 0.00 $\pm$ 0.00 12.49 $\pm$ 0.00
32 44.59 $\pm$ 0.03 44.59 $\pm$ 0.03 0.02 $\pm$ 0.00 9.21 $\pm$ 0.00 0.00 $\pm$ 0.00 12.49 $\pm$ 0.00
48 46.71 $\pm$ 3.20 46.71 $\pm$ 3.20 0.02 $\pm$ 0.00 9.21 $\pm$ 0.00 0.00 $\pm$ 0.00 12.49 $\pm$ 0.00
64 48.67 $\pm$ 0.13 48.67 $\pm$ 0.13 0.02 $\pm$ 0.00 9.21 $\pm$ 0.00 0.00 $\pm$ 0.00 12.49 $\pm$ 0.00
Character Recognition
8 48.76 $\pm$ 0.00 48.76 $\pm$ 0.00 0.00 $\pm$ 0.00 0.32 $\pm$ 0.00 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01
16 48.62 $\pm$ 0.23 48.62 $\pm$ 0.23 0.00 $\pm$ 0.00 0.14 $\pm$ 0.00 0.00 $\pm$ 0.00 0.05 $\pm$ 0.00
32 48.98 $\pm$ 0.08 48.98 $\pm$ 0.08 0.00 $\pm$ 0.00 0.14 $\pm$ 0.00 0.00 $\pm$ 0.00 0.05 $\pm$ 0.00
48 48.91 $\pm$ 0.09 48.91 $\pm$ 0.09 0.00 $\pm$ 0.00 0.14 $\pm$ 0.00 0.00 $\pm$ 0.00 0.05 $\pm$ 0.00
64 48.72 $\pm$ 0.06 48.72 $\pm$ 0.06 0.00 $\pm$ 0.00 0.14 $\pm$ 0.00 0.00 $\pm$ 0.00 0.05 $\pm$ 0.00
Results. Table 17 presents the performance of LoRA-based adaptation across a range of rank values ($r\in\{8,16,32,48,64\}$) for both Action Recognition (AR) and Character Recognition (CR) tasks, across all prompt formats. We report exact match (EM) and ROUGE-F1 averaged over three runs. Increasing the rank beyond $r=8$ yields no consistent improvements across tasks or formats. Performance on binary prompts remains close to random, while performance on multiple-choice and open-ended formats stays near zero across all ranks. These results suggest that LoRA, even with increased capacity, is insufficient for capturing the fine-grained temporal and semantic dependencies required by our evaluation protocol. Given the lack of benefit from increasing rank—and the added parameter cost—it is inefficient to scale LoRA rank beyond $r=8$. Accordingly, all results reported in the main paper use $r=8$, while extended comparisons with higher ranks are presented here for completeness.

### v2

@@
 Yet, their use in fine-grained, temporally sensitive evaluation tasks remains limited and requires targeted adaptation.
-We introduce a evaluation protocol targeting two recognition tasks – action recognition and character recognition – each assessed across binary, multiple-choice, and open-ended formats.
-To support this, we present universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a method for adapting VLMs to rollout evaluation under data and compute constraints.
-We conduct a large-scale study comparing full, partial, and parameter-efficient finetuning across task formats, context lengths, sampling strategies, and data compositions.
-The resulting unified evaluator matches the performance of task-specific baselines using a single checkpoint.
-Human studies confirm strong alignment with human judgments, establishing universe as a scalable, semantics-aware evaluator for world models.
+We introduce an evaluation protocol targeting two recognition tasks – action recognition and character recognition – each assessed across binary, multiple-choice, and open-ended formats.
+To support this, we present universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a VLM-based evaluator for video world model rollouts adapted under data and compute constraints.
+In our extensive experiments totaling over 5,154 GPU-days, we explore full, partial, and parameter-efficient adaptation methods across various task formats, context lengths, sampling methods, and data compositions.
+The resulting unified evaluator achieves parity with task-specific checkpoints.
+Human studies across seven diverse environments confirm strong alignment with human judgments, establishing universe as a lightweight, adaptable, and semantics-aware evaluator for video world models.
 1 Introduction
-World models are generative models trained to predict future observations conditioned on past observations and actions [37, 41, 5].
-They offer a powerful abstraction for learning, reasoning, and planning in complex interactive environments, and are rapidly becoming foundational in domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77].
-As their capabilities grow, a persistent challenge remains: evaluation.
-Rollouts produced by world models are visually rich, temporally grounded, and semantically structured.
-Evaluating them requires: (i) precise alignment between generated frames and control sequences at the timestamp level [106], and (ii) consistent tracking of entities and their semantics over time [56].
-Existing evaluation protocols offer limited insight in this regime.
-Traditional visual generation metrics are ill-suited to the structured and dynamic nature of rollouts: (i) early distributional metrics focus on static images and are sensitive to low-level variations [85, 46, 12], (ii) motion-aware metrics such as FVD [97] capture temporal patterns but lack semantic grounding, and (iii) multimodal metrics improve semantic alignment but ignore action conditioning and timestamp-level fidelity [52].
-While human evaluation remains the gold standard [2, 6], it is expensive and difficult to scale.
-Similarly, emerging text-to-video benchmarks [66, 50, 62] focus on open-ended generation and neglect the fine-grained control and temporal coherence essential to evaluating world model outputs.
-[…]
-VLM have demonstrated strong generalization across multimodal tasks by integrating visual and linguistic reasoning [61, 30, 24, 99, 1, 64, 26, 72], and are increasingly used to evaluate generative models [59, 70, 63, 20].
-We extend this direction, hypothesizing that VLMs can serve as fine-grained evaluators of world model rollouts, capturing semantic consistency, temporal coherence, and alignment with control inputs.
-Virtual environments offer a favorable testbed for this purpose, as they expose full access to timestamped actions, object states, and agent behaviors, enabling controlled and systematic evaluation.
-Figure 1: Performance and efficiency of universe (orange bars throughout) compared to task-specific baselines (multiple colours).
+World models – generative models trained to predict future observations conditioned on past observations and actions (Ha & Schmidhuber, 2018; Hafner et al., 2025; Alonso et al., 2024) – are rapidly becoming central to interactive AI.
+They provide a powerful abstraction for learning, reasoning, and planning in complex interactive environments, and underpin advances in neural game engines (Kanervisto et al., 2025; Guo et al., 2025; Gao et al., 2025; Chen et al., 2025), embodied AI (Du et al., ; Yang et al., 2024), and autonomous driving (Russell et al., 2025; Hu et al., 2023a; Ni et al., 2025).
+Yet, evaluating world models remains a bottleneck.
+Rollouts are semantically rich and temporally grounded, requiring metrics that assess (i) alignment between generated frames and action sequences at the timestamp level (Yang et al., 2024), and (ii) consistent entity tracking over time (Kanervisto et al., 2025).
+Existing approaches fall short: (i) early distributional metrics focus on images and are sensitive to low-level variations (Salimans et al., 2016; Heusel et al., 2017; Binkowski et al., 2018), (ii) motion-aware metrics like FVD (Unterthiner et al., 2018) lack semantic grounding, and (iii) multimodal metrics ignore timestamp-level action conditioning (Jayasumana et al., 2024).
+Emerging text-to-video benchmarks (Liu et al., 2024b; Huang et al., 2024; Liao et al., 2024) focus on open-ended generation but neglect the fine-grained control central to world model evaluation.
+Even cutting-edge LLMs fail in this setting (Appendix G.1, Figure 15).
+Human evaluation remains the gold standard (Agarwal et al., 2025; Analysis, 2024), however, it remains costly and hard to scale.
+To address this gap, we propose a novel evaluation protocol targeting two important dimensions of rollout quality: action alignment and character consistency, formalized as recognition tasks – Action Recognition (AR), Character Recognition (CR) – across formats of varying complexity.
+The protocol provides a foundation for semantic evaluation of rollouts, and extends upon existing evaluation method by providing insight into semantic quality of generated rollouts.
+Inspired by the LLM-as-a-Judge research direction Zheng et al.
+(2023), we explore whether Vision-Language Models can serve as effective evaluators in this setting.
+VLMs have demonstrated strong generalization across multimodal tasks (Li et al., 2023; Driess et al., 2023; Chen et al., 2023; Wang et al., 2024; Abdin et al., 2024; Liu et al., 2024a; Deitke et al., 2024; McKinzie et al., 2024), and show promise as automatic judges of generative models (Lee et al., 2024; Mañas et al., 2024; Lin et al., 2024; Chen et al., 2024).
+Yet, off-the-shelf VLMs struggle with the temporal grounding and domain-specific knowledge (see Sec.
+4, Zero-Shot Evaluation).
+Limited resources and sparse text supervision introduce another layer of complexity to the setting.
+We therefore conduct a systematic study of adaptation strategies under realistic data and compute constraints.
+Throughout the study totaling over 5,154 GPU-days, we analyze the impact of supervision regime, frame sampling strategy, visual context length, and training budget.
+The result of the study is universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a lightweight, adaptable, and semantics-aware evaluator for world models.
+universe achieves parity with six task-specific checkpoints while using a single unified model (see Figure 1.
+To validate reliability, we conduct a large-scale human study on rollouts spanning diverse environments, model scales, and rollout fidelities.
+universe’s judgments show strong agreement with human ratings, demonstrating its effectiveness as a practical, semantics-aware evaluator—particularly valuable in settings where ground-truth annotations are unavailable or prohibitively expensive.
+Figure 1: Performance and efficiency of universe (orange bars throughout) compared to task-specific baselines (multiple colours), all models trained for 10 epochs.
 Left and Center: Action recognition and Character Recognition accuracy across binary, multiple-choice, and open-ended settings.
 Right: Sample efficiency – our adaptation recipe achieves strong performance with substantially fewer training samples per epoch.
-universe matches the performance of task-specific baselines while using a single checkpoint, enabled by mixed supervision, efficient frame sampling, and lightweight partial fine-tuning.
-[…]
-[…]
-To support this protocol, we introduce universe (UNIfied Vision-language Evaluator for Rollouts in Simulated Environments), a method for adapting VLMs to structured rollout evaluation.
-universe emerges from a systematic study of adaptation strategies under realistic data and compute constraints.
-We analyze the impact of supervision regime, frame sampling strategy, visual context length, and training budget, arriving at an adaptation recipe that combines mixed supervision, efficient frame selection, and lightweight fine-tuning.
-To assess the reliability of universe, we conduct a human annotation study using rollouts generated by WHAM [56], a publicly available world model trained in a complex interactive environment.
-We evaluate the rollouts using universe and collect human judgments on the same data to compare responses.
-We find that universe’s judgments exhibit strong alignment with human ratings, confirming its potential for automated evaluation world model rollouts, particularly when explicit ground truth is unavailable or costly to obtain.
-In summary, this work makes three key contributions:
-  1. I
-    We introduce a structured protocol for evaluating world model rollouts, based on action and character recognition tasks with increasing levels of complexity;
-  2. II
-[…]
-  3. III
-    We show that universe matches the performance of task-specific baselines using a single checkpoint, demonstrating its ability to generalize across tasks with a unified model (Figure 1).
-Furthermore, its predictions align closely with human judgments, validating its effectiveness as a scalable, semantics-aware evaluator for world model rollouts.
 2 Related Work
 Challenges in Evaluating World Models.
-World models are generative systems that learn predictive representations of environment dynamics [37], originally proposed for model-based RL [92] and now central to domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77].
-Recent models such as Dreamer v1–3 [39, 38, 40, 41], MuZero [86], IRIS [74], UniSim [106], and DIAMOND [5] have improved rollout fidelity and controllability.
-Yet evaluation largely focuses on downstream success metrics—e.g., game score or goal completion [9, 54, 36, 7, 8]—which provide only coarse, indirect signals of rollout quality.
-Genie [17, 78] decouples world model learning and agent training, but its evaluation still emphasizes visual quality and control, without probing semantic or causal fidelity.
-Cosmos [2] proposes a structured protocol that combines FID/FVD with structure-from-motion-based 3D consistency checks and human ratings on instruction following, object permanence, and visual verity.
+World models are generative systems that learn predictive representations of environment dynamics (Ha & Schmidhuber, 2018), originally proposed for model-based RL (Sutton, 1991) and now central to domains such as neural game engines (Kanervisto et al., 2025; Guo et al., 2025; Gao et al., 2025; Chen et al., 2025), embodied AI (Du et al., ; Yang et al., 2024), and autonomous driving (Russell et al., 2025; Hu et al., 2023a; Ni et al., 2025).
+Recent models such as Dreamer v1–3 (Hafner et al., 2020; ; 2023; 2025), MuZero (Schrittwieser et al., 2020), IRIS (Micheli et al., 2023), UniSim (Yang et al., 2024), and DIAMOND (Alonso et al., 2024) have improved rollout fidelity and controllability.
+Yet evaluation largely focuses on downstream success metrics—e.g., game score or goal completion (Bellemare et al., 2013; Kaiser et al., 2020; Guss et al., 2021; Baker et al., 2022; Beattie et al., 2016)—which provide only coarse, indirect signals of rollout quality.
+Genie (Bruce et al., 2024; Parker-Holder et al., 2024) decouples world model learning and agent training, but its evaluation still emphasizes visual quality and control, without probing semantic or causal fidelity.
+Cosmos (Agarwal et al., 2025) proposes a structured protocol that combines FID/FVD with structure-from-motion-based 3D consistency checks and human ratings on instruction following, object permanence, and visual verity.
 While insightful, this approach is tied to simulator-specific infrastructure and requires costly manual comparison.
-Human-in-the-loop protocols such as the Video Generation Arena [6] also rely on pairwise comparison to assess rollout quality.
+Human-in-the-loop protocols such as the Video Generation Arena (Analysis, 2024) also rely on pairwise comparison to assess rollout quality.
 These methods, though informative, are expensive and hard to scale.
 Evaluation Metrics and Protocols for Visual Generation.
-Early evaluations of generative models relied on full-reference metrics such as PSNR and SSIM [100], which capture pixel-level and perceptual similarity but are sensitive to spatial misalignments and fail to reflect semantic fidelity.
-To address this, distributional metrics like Inception Score (IS) [85], Fréchet Inception Distance (FID) [46], and Kernel Inception Distance (KID) [12].
-Other proposals such as PPL [57], Parzen likelihoods [34], and HYPE [110] attempt to quantify perceptual smoothness or human realism, but remain focused on static images.
-For video generation, FVD [97] generalizes FID using I3D features [18], introducing a motion-aware distributional baseline.
+Early evaluations of generative models relied on full-reference metrics such as PSNR and SSIM (Wang et al., 2004), which capture pixel-level and perceptual similarity but are sensitive to spatial misalignments and fail to reflect semantic fidelity.
+To address this, distributional metrics like Inception Score (IS) (Salimans et al., 2016), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Kernel Inception Distance (KID) (Binkowski et al., 2018).
+Other proposals such as PPL (Karras et al., 2019), Parzen likelihoods (Goodfellow et al., 2014), and HYPE (Zhou et al., 2019) attempt to quantify perceptual smoothness or human realism, but remain focused on static images.
+For video generation, FVD (Unterthiner et al., 2018) generalizes FID using I3D features (Carreira & Zisserman, 2017), introducing a motion-aware distributional baseline.
 Yet, FVD also lacks semantic grounding and does not account for causal structure or goal alignment.
 To improve semantic grounding, metrics based on text-image alignment have been proposed.
-CLIPScore [45] and CLIPSIM [102] compute similarity between generated visuals and textual or visual references using CLIP embeddings [82], while Jayasumana et al.
-[52] extend this to distributional comparisons via MMD.
+CLIPScore (Hessel et al., 2021) and CLIPSIM (Wu et al., 2021) compute similarity between generated visuals and textual or visual references using CLIP embeddings (Radford et al., 2021), while Jayasumana et al.
+(2024) extend this to distributional comparisons via MMD.
 However, all operate at the frame level.
 Structured evaluation protocols using vision-language reasoning have also emerged.
-VQA Accuracy [70] uses LLMs to score answers on static image questions, and VQAScore [63] probes alignment via templated binary queries.
+VQA Accuracy (Mañas et al., 2024) uses LLMs to score answers on static image questions, and VQAScore (Lin et al., 2024) probes alignment via templated binary queries.
 Lee et al.
-[59] propose VLM evaluator to evaluate other VLMs responses given user criteria.
+(2024) propose VLM evaluator to evaluate other VLMs responses given user criteria.
 These approaches introduce task structure but remain limited to single-frame evaluation.
-Recent text-to-video (T2V) benchmarks such as EvalCrafter [66], VBench [50], and DEVIL [62] introduce curated prompts and metrics covering text alignment, motion realism, and perceptual quality.
+Recent text-to-video (T2V) benchmarks such as EvalCrafter (Liu et al., 2024b), VBench (Huang et al., 2024), and DEVIL (Liao et al., 2024) introduce curated prompts and metrics covering text alignment, motion realism, and perceptual quality.
 While these protocols push forward evaluation of open-ended video generation, they lack timestamp-level action grounding.
 Vision-Language Model Adaptation.
-VLMs have emerged as powerful tools for multimodal understanding, demonstrating strong performance across tasks such as captioning, retrieval, visual question answering, and instruction following [61, 30, 24, 99, 1, 64, 26, 72].
+VLMs have emerged as powerful tools for multimodal understanding, demonstrating strong performance across tasks such as captioning, retrieval, visual question answering, and instruction following (Hendriksen et al., 2022; Li et al., 2023; Driess et al., 2023; Chen et al., 2023; Wang et al., 2024; Abdin et al., 2024; Liu et al., 2024a; Deitke et al., 2024; McKinzie et al., 2024).
 Adaptation approaches can be broadly categorized into prompt-level and weight-level methods.
-One prominent prompt-level adaptation techniques is prompt tuning, which injects task information directly into the input space [75, 111, 103], and in-context learning (ICL), where models such as GPT-3 [16] and Flamingo [4] condition on task demonstrations at inference time without updating parameters.
-Retrieval-augmented generation (RAG) [60] combines parametric models with non-parametric memory, and multimodal variants incorporate external visual or auditory context [49, 22].
+One prominent prompt-level adaptation techniques is prompt tuning, which injects task information directly into the input space (Miyai et al., 2023; Zhou et al., 2024; Wu et al., 2024a), and in-context learning (ICL), where models such as GPT-3 (Brown et al., 2020) and Flamingo (Alayrac et al., 2022) condition on task demonstrations at inference time without updating parameters.
+Retrieval-augmented generation (RAG) (Lewis et al., 2020) combines parametric models with non-parametric memory, and multimodal variants incorporate external visual or auditory context (Hu et al., 2023b; Chen et al., 2022a).
 While lightweight, these approaches are limited in their ability to model temporal dependencies or align with structured rollouts.
 Weight-level adaptation enables stronger domain alignment but incurs higher computational cost.
-Full finetuning remains effective yet costly, while partial finetuning [107] offers a trade-off by updating only selected layers.
-Parameter-efficient finetuning (PEFT) provides a scalable alternative and can be grouped into low-rank and adapter-based strategies [42].
-Low-rank methods, such as LoRA [48], inject rank-constrained updates into frozen layers.
-Recent extensions improve upon this via weight decomposition [65], quantization-aware adaptation [27, 105], mixture-of-experts routing [104], and long-context support [25].
-Adapter-based methods insert lightweight modules between frozen layers to enable modular adaptation with minimal overhead [69, 109].
+Full finetuning remains effective yet costly, while partial finetuning (Ye et al., 2023) offers a trade-off by updating only selected layers.
+Parameter-efficient finetuning (PEFT) provides a scalable alternative and can be grouped into low-rank and adapter-based strategies (Han et al., 2024).
+Low-rank methods, such as LoRA (Hu et al., 2022), inject rank-constrained updates into frozen layers.
+Recent extensions improve upon this via weight decomposition (Liu et al., ), quantization-aware adaptation (Dettmers et al., 2023; Xu et al., 2024), mixture-of-experts routing (Wu et al., 2024b), and long-context support (Chen et al., ).
+Adapter-based methods insert lightweight modules between frozen layers to enable modular adaptation with minimal overhead (Luo et al., 2023; Zhao et al., 2024).
 A parallel line of work investigates multimodal few-shot learning.
-Frozen [95] was among the first to explore this setting, followed by works combining prompting and ICL for improved sample efficiency [53, 89], and works introducing a learnable meta-mapper to bridge frozen VLM components for few-shot meta-learning [76].
-[…]
+Frozen (Tsimpoukelli et al., 2021) was among the first to explore this setting, followed by works combining prompting and ICL for improved sample efficiency (Jin et al., 2022; Song et al., 2022), and works introducing a learnable meta-mapper to bridge frozen VLM components for few-shot meta-learning (Najdenkoska et al., 2023).
+Our Focus.
+While prior efforts have explored related challenges, none directly address the evaluation of the structured, action-conditioned fidelity and semantics of world model rollouts using adapted VLM.
+To this end, we introduce: (i) an evaluation protocol for world model rollouts, targeting fine-grained, temporally grounded assessment of semantic fidelity; (ii) universe, a VLM-based method to support the protocol.
+We validate its alignment with human judgments and demonstrate its scalability and semantic sensitivity across rollout conditions.
 3 Methodology
@@
 Evaluation quality is measured by comparing $\hat{A}$ to the reference answer $A$ using semantic similarity metrics.
-[…]
+Evaluation Protocol.
+We define two structured recognition tasks: (i) Action Recognition (AR): Assesses whether generated sequences accurately reflect the effects of agent actions given the segment; (ii) Character Recognition (CR): Evaluates whether entities maintain consistent identity and appearance across time.
+Each task is framed as a visual QA problem: the evaluator receives a sequence of frames and a natural language prompt (binary, multiple-choice, or open-ended), and generates a textual response.
+Outputs are scored using Exact Match (EM) and ROUGE-F1 (ROUGE), capturing both literal and semantic alignment with the reference answer.
+Metric details are in Appendix E.2.
 Dataset Construction.
 Effective VLM adaptation for rollout evaluation requires a dataset that (i) captures realistic human behavior in interactive environments, and (ii) aligns with prior work in simulated settings to support comparability and reproducibility.
-Additionally, since WHAM – the world model used in our evaluation – is trained on Bleeding Edge, it is critical that the adaptation dataset match its distribution.
-To satisfy these constraints, we partnered with Ninja Theory and curated a dataset from both internal and public Bleeding Edge gameplay recordings, focusing on the Skygarden map used in WHAM [56].
-This dataset provides high visual and behavioral diversity [79], includes a publicly available evaluation split, and is closely aligned with prior work in the domain [56, 79, 94, 88, 28], enabling cross-method comparison.
-Data preparation proceeds in three stages: (i) Preprocessing:Segment gameplay into 14-frame clips with synchronized video, control logs, and metadata; (ii) Description Generation:Convert structured annotations (e.g., actions, agent states) into natural language summaries; (iii) Question-Answer Pair Construction:Generate six QA pairs per clip (binary, multiple-choice, and open-ended) spanning both AR and CR tasks.
+To satisfy these constraints, we partnered with Ninja Theory and curated a dataset from both internal and public Bleeding Edge gameplay recordings, focusing on the Skygarden evnironment (Kanervisto et al., 2025).
+This dataset provides high visual and behavioral diversity (Pearce et al., 2025), includes a publicly available evaluation split, and is closely aligned with prior work in the domain (Kanervisto et al., 2025; Pearce et al., 2025; Tot et al., 2025; Sharma et al., 2024; Devlin et al., 2021), enabling fair comparison.
+Data preparation proceeds in three stages: (i) Preprocessing:Segment gameplay into 14-frame clips with synchronized video, control logs, and metadata; (ii) Description Generation:Convert structured annotations (e.g., actions, agent states) into natural language summaries; (iii) Question-Answer Pair Construction:Generate six complimentary QA pairs per clip (binary, multiple-choice, and open-ended) spanning both AR and CR tasks.
 The final dataset contains 32.453 training clips and 8.113 validation clips, yielding 194.718 and 48.678 QA pairs, respectively.
-See Appendix C for details.
+See Appendix D for details.
 Model Architecture.
-We adapt a model from the PaliGemma family [10, 91], consisting of a vision encoder $\mathcal{M}_{V}$, a projection head $\mathcal{M}_{P}$, and a language decoder $\mathcal{M}_{L}$.
-Based on initial zero-shot evaluations (Appendix F.1), we use a single configuration for all experiments—PaliGemma 2 3b, which includes a 2B-parameter Gemma 2 decoder pretrained on 2T tokens.
+We adapt a model from the PaliGemma family (Beyer et al., 2024; Steiner et al., 2024), consisting of a vision encoder $\mathcal{M}_{V}$, a projection head $\mathcal{M}_{P}$, and a language decoder $\mathcal{M}_{L}$.
+Based on initial zero-shot evaluations (Appendix G.2), we use a single configuration for all experiments—PaliGemma 2 3b, which includes a 2B-parameter Gemma 2 decoder pretrained on 2T tokens.
 Input frames are resized to $224\times 224$ and tokenized into 256 patches each.
-Model achitecture details are in Appendix D.1.
+Model achitecture details are in Appendix E.1.
 Each model input sequence $S=\{S_{\mathcal{I}},S_{\mathcal{T}}^{\text{PREF}},S_{\mathcal{T}}^{\text{SUFF}}\}$ consists of: visual tokens $S_{\mathcal{I}}$ from $k$ frames, a textual prefix $S_{\mathcal{T}}^{\text{PREF}}$ containing the task-language cue and question, and a suffix $S_{\mathcal{T}}^{\text{SUFF}}$ with the expected answer (used only during training).
 This format allows the decoder to attend jointly over visual and textual context.
-Full prompt details are provided in Appendix D.1.
+Full prompt details are provided in Appendix E.1.
 Training Objective. We optimize a causal language modeling loss on the answer suffix:
@@
 (iv) Single-component fine-tuning:Only one module—vision, projection, or language—is updated.
-(v) Parameter-efficient fine-tuning:We apply LoRA [48] adapters to attention and MLP layers in vision and language components: $\mathbf{W}\leftarrow\mathbf{W}+\frac{\alpha}{r}\mathbf{A}\mathbf{B},\quad\alpha=8,\;r\in\{8,16,32,48,64\}$.
+(v) Parameter-efficient fine-tuning:We apply LoRA (Hu et al., 2022) adapters to attention and MLP layers in vision and language components: $\mathbf{W}\leftarrow\mathbf{W}+\frac{\alpha}{r}\mathbf{A}\mathbf{B},\quad\alpha=8,\;r\in\{8,16,32,48,64\}$.
 Frame Sampling Policy.
@@
 Designed for constrained compute and limited supervision, universe delivers strong generalization across our evaluation protocol using a single, partially tuned model.
-The method combines three key components:
-  1. I
-    Partial fine-tuning: We update only the projection head ($\theta_{P}$), training just 0.07% of model parameters.
-Despite this minimal footprint, it achieves the second-best performance among all strategies—trailing only vision encoder tuning, which requires $\sim$11of parameters and incurs significantly higher compute cost.
-  2. II
-    Efficient frame sampling: Each input sequence includes $k=8$ frames sampled uniformly from a 14-frame rollout.
+The method combines three main components: I Partial fine-tuning:We update only the projection head ($\theta_{P}$), training just 0.07% of model parameters.
+Despite this minimal footprint, it achieves the second-best performance among all strategies—trailing only vision encoder tuning, which requires $\sim$11% of parameters and incurs significantly higher compute cost.
+II Efficient frame sampling:Each input sequence includes $k=8$ frames sampled uniformly from a 14-frame rollout.
 This sparsity-aware strategy maintains long-range temporal structure while reducing token count and enabling efficient batching.
-  3. III
-    Mixed supervision: We train on a hierarchical mixture of tasks and QA formats.
-The task distribution favors Action Recognition ($\alpha_{\texttt{AR}}=0.8$) due to its stronger causal grounding.
+III Mixed supervision:We train on a hierarchical mixture of tasks and QA formats.
+The task distribution favors Action Recognition ($\alpha_{\texttt{AR}}=0.8$) as it converges slower.
 Within each task, we emphasize open-ended questions ($\beta_{\texttt{OE}}=0.8$), while maintaining smaller proportions of binary ($\beta_{\texttt{binary}}=0.15$) and multiple-choice ($\beta_{\texttt{MC}}=0.05$) examples.
 4 Experiments
-[…]
-Figure 2: Comparison of universe and baseline models on Action and Character Recognition.
+We evaluate universe on ground truth video data, focusing on AR and CR across binary, multiple-choice, and open-ended formats.
+Our goals are twofold: (i) to benchmark performance against zero-shot and fine-tuned baselines, and (ii) to assess the trade-offs between adaptation strategies under constrained supervision and compute.
+Figure 2: Comparison of universe and baseline models on Action and Character Recognition, all models trained for 1 epoch.
 Left: universe outperforms all baselines on AR.
 Right: On CR, it ranks third, behind models with either full vision encoder tuning or task-specific training with greater supervision.
-Trained under a unified protocol with minimal parameter updates (0.07%) and reduced per-task data, universe delivers strong performance across both tasks, highlighting its efficiency and generalization.
-Baselines.
-We compare universe against two classes of baselines: (i) Zero-shot VLMs:Seven off-the-shelf models, including VideoLLaMA3 (2B, 7B) [14] and three PaliGemma models: version 1 (3B) and version 2 (3B and 10B) [10, 91], evaluated without domain adaptation using an 8-frame visual context window.
-[^1] We also experimented with CLIPScore-based evaluation (Appendix F.2); results underperformed relative to selected baselines and were constrained to predefined candidate sets, further underscoring the need for model adaptation.
-(ii) Fine-tuned PaliGemma 2:Variants adapted via full, partial, and parameter-efficient tuning.
-This backbone is selected based on a sweep over PaliGemma variants, using zero-shot performance as a guide (Appendix F.1).
-The adaptation space includes 8 primary baselines: (i) Single-component fine-tuning: tuning only the vision encoder ($\mathcal{F}_{\text{V}}$), the multimodal projector ($\mathcal{F}_{\text{P}}$), or the language head ($\mathcal{F}_{\text{L}}$); (ii) Two-component fine-tuning: jointly tuning pairs of components—$\mathcal{F}_{\text{V+P}}$, $\mathcal{F}_{\text{V+L}}$, and $\mathcal{F}_{\text{P+L}}$; (iii) Full-model fine-tuning: tuning all components simultaneously ($\mathcal{F}_{\text{V+P+L}}$); (iv) LoRA-based tuning: Parameter-efficient adaptation with rank $r=8$, selected after observing minimal performance variation across $r\in\{8,16,32,48,64\}$ (see Appendix F.3 for details).
-All models are trained using 8-frame clips and a single epoch.
-Results.
-Figure 2 (left, center) summarizes performance across Action Recognition (AR) and Character Recognition (CR).
-Zero-shot models perform poorly across both tasks: VideoLLaMA3 variants fall below 12.7% on AR and 6.4% on CR; PaliGemma variants reach up to 29.7% on AR and 17.2% on CR.
-These results confirm that general-purpose VLMs lack the temporal grounding and domain-specific semantics required for structured rollout evaluation.
-In contrast, universe achieves strong performance—outperforming all models on AR and ranking third on CR.
-The top two CR models either fine-tune the full vision encoder, requiring updates to $\sim$400M parameters, or fine-tune the multimodal projector with 5 $\times$ more CR supervision, each trained in isolation on a single task and prompt format.
-universe, by comparison, tunes only the 2.66M-parameter projector (0.07% of the model), under a unified protocol spanning both tasks, all prompt formats, and reduced per-task supervision.
-Its performance under these constraints underscores the efficiency and generality of our adaptation strategy for temporally grounded evaluation.
+Trained under a unified protocol with minimal parameter updates (0.07%) and reduced per-task data, universe delivers strong performance across both tasks.
+[…]
+[…]
 5 Analysis
-[…]
+We find a consistent performance gap between AR and CR, highlighting the greater temporal and causal complexity of action understanding.
+This motivates our focus on AR as the more challenging and diagnostic task.
+Below, we analyze how adaptation choices shape performance on AR.
 Supervision and Temporal Context.
@@
 By independently and jointly varying the number of training epochs and input frames, we disentangle the contributions of model capacity and temporal context to task success.
-Figure 3: Exact Match accuracy over training epochs for Action Recognition (AR, top) and Character Recognition (CR, bottom).
-AR improves gradually with supervision, while CR converges to high accuracy with minimal training.
-[…]
+Results.
+CR converges rapidly, achieving over 97% exact match after 12.5% of an epoch (~4K samples; Figure 5, bottom), and shows minor improvement with further training.
+In contrast, AR improves only modestly under extended training when limited to a single frame (Figure 5, top), suggesting that supervision alone is insufficient in the absence of temporal information.
+Motivated by this, we jointly scale both supervision and input length, varying the number of frames and epochs.
+As shown in Figure 3, performance on AR improves consistently across all formats, with the best results achieved under combined scaling.
 Figure 3: Action Recognition performance as a function of training supervision (epochs) and temporal context (number of frames), evaluated across all formats.
 Performance improves along both axes, with highest accuracy achieved when both dimensions are scaled.
-[…]
+Figure 4: Effect of frame sampling strategy on Action Recognition performance across all formats.
+Uniform-$n$ sampling (orange) consistently outperforms first-$n$ (blue), with especially large gains at low frame counts, and maintains an advantage as temporal context increases.
+Temporal Sampling Strategies.
+Following the observation that AR requires both extended supervision and temporally rich input, we examine how frame selection impacts performance.
+We compare first-$n$ sampling, which selects the first $n$ consecutive frames from each rollout, to uniform-$n$ sampling, which draws $n$ evenly spaced frames across the entire sequence.
+We conduct experiments at varying context lengths, using $n\in\{1,2,\dots,8\}$ frames.
+to evaluate the impact of both sampling method and input horizon.
+Figure 5: Exact Match accuracy for Action Recognition and Character Recognition.
 Results.
@@
 Gains persist even at 8 frames, where uniform sampling maintains an advantage across formats.
-Figure 5: Effect of frame sampling strategy on Action Recognition performance across all formats.
-Uniform-$n$ sampling (orange) consistently outperforms first-$n$ (blue), with especially large gains at low frame counts, and maintains an advantage as temporal context increases.
-[…]
+Optimizing Data Mix for Unified Multi-Task Evaluation.
+We analyze how training data composition affects multi-task performance in universe, with the goal of enabling a single model to generalize across AR and CR.
+Specifically, we study how the task-level ratio $\alpha$ and format-level ratio $\beta$ influence performance across evaluation settings.
+We first conduct a hierarchical ablation to identify an optimized data mixture, then assess its impact by comparing against a default task mix with uniform sampling.
 Figure 6: Hierarchical ablation of training data composition for universe.
@@
 6 Evaluating World Model Rollouts with universe
-We evaluate universe as an automated evaluator of world model rollouts using the WHAM benchmark [56], which provides pretrained world models and an evaluation set.
-Our analysis focuses on two axes: (i) in-domain accuracy, measured on Skygarden—the environment used during fine-tuning—and (ii) generalization to six unseen environments.
-We compare universe’s predictions on samples from two world models: (i) WHAM-140M, trained on Skygarden with lower-quality rollouts ($128\times 128$ resolution), and (ii) WHAM-1.6B, trained on a diverse environment suite with higher-resolution output ($300\times 180$).
-We prepare 30 rollouts for each model-environment pair, yielding a total of 240 rollouts.
-Each rollout is segmented into 14-frame clips and paired with six natural language questions, following our evaluation protocol.
-universe answers each question using majority voting over five greedy decoding samples.
-Human annotators then rate each response on a four-point ordinal scale: Correct, Partially Correct, Incorrect, and Unclear.
-Each response is independently rated by two annotators; in cases of disagreement, a third annotator serves as adjudicator.
-Inter-annotator agreement is quantified using Cohen’s $\kappa$.
-Full details of the annotation protocol are provided in Appendix E.
+We evaluate the reliability of universe through a human study spanning eight distinct settings that vary in model scale, training data diversity, and output resolution.
+Our analysis considers two axes: (i) in-domain accuracy, measured on Skygarden, and (ii) generalization to six previously unseen environments.
+Concretely, we study rollouts generated by: (i) a large-scale model trained across multiple environments with higher-resolution rollouts ($300\times 180$), and (ii) a smaller model trained on a single environment with lower-resolution rollouts ($128\times 128$).
+These two model families expose complementary challenges: resolution mismatch, domain coverage, and rollout fidelity.
+We construct eight evaluation settings: settings 1–7 draw from the large-scale model across diverse environments, while setting 8 uses the smaller model in the fine-tuning environment.
+Each setting contains 30 rollouts, paired with six natural-language questions from our evaluation protocol, yielding 240 rollouts in total.
+universe answers each question via majority vote over five greedy decoding samples.
+Human annotators rate responses on a four-point ordinal scale (Correct, Partially Correct, Incorrect, Unclear), with double annotation and adjudication on disagreements.
+Inter-rater reliability is measured using Cohen’s $\kappa$.
+Full details of the annotation protocol are provided in Appendix F.
 Results.
-Figure 8 summarizes graded accuracy across models and environments.
-We observe a significant performance gap between rollouts from WHAM-140M and WHAM-1.6B.
-Despite being in-domain, WHAM-140M yields lower evaluation accuracy likley due to a resolution mismatch: its $128\times 128$ frames are upsampled to $224\times 224$ for universe input.
-In contrast, WHAM-1.6B produces higher-quality inputs, enabling more reliable evaluation.
-On WHAM-1.6B Skygarden rollouts, universe achieves 75.02% graded accuracy on AR and 90.00% on CR.
-When applied to six previously unseen environments, AR accuracy remains relatively stable with the lowest accuracy of 66.11% for environment G, and highest accuracy of 83.33% for environment E.
-Overall, the results indicate strong generalization, especially for identity-focused recognition tasks.
-Cohen’s $\kappa$ scores further reflect the interpretability of universe’s outputs: agreement is substantial overall ($\kappa=0.73$), with the lowest in Environment G ($\kappa=0.59$) and highest in Environment A on WHAM-1.6B rollouts ($\kappa=0.91$).
-These results highlight that universe remains aligned with human judgments.
-Figure 8: Graded accuracy of universe across rollouts from WHAM-140M (Env.
-A) and WHAM-1.6B (Envs.
-A–G).
-Performance improves markedly with higher-fidelity rollouts from WHAM-1.6B, even in out-of-domain settings (B–G).
-Cohen’s $\kappa$ (above bars) reflects inter-rater agreement.
+Figure 9 reports graded accuracy across all settings.
+Rollouts from the smaller, single-environment model yield lower evaluation accuracy, likely due to resolution mismatch, while the larger, multi-environment model provides higher-quality inputs.
+These results demonstrate that universe generalizes across model scales, rollout fidelities, and environments, while remaining closely aligned with human judgments.
+Figure 8: Example frames from the eight evaluation settings, spanning different model scales, rollout fidelities, and environments.
+Note the resolution difference: Settings 1-7 feature $300\times 180$ rollouts, whereas Setting 8 uses $128\times 128$.
+Figure 9: Graded accuracy of universe across the eight evaluation settings.
+Performance improves with higher-fidelity rollouts and remains stable across unseen environments (2–7).
+Cohen’s $\kappa$ indicates substantial inter-rater agreement.
 7 Conclusion
@@
 To support this, we propose universe, a unified method for adapting VLMs to this setting through mixed supervision, efficient frame sampling, and lightweight fine-tuning.
-Our large-scale study demonstrates that universe matches the performance of task-specific baselines using a single checkpoint and aligns closely with human judgments, establishing it as a scalable, semantics-aware evaluator for evaluating world models, particularly when explicit ground truth is unavailable or costly to obtain.
-Limitations While universe performs well in simulation, its generalization beyond simulated environments remains an open challenge.
-The evaluation protocol targets two fidelity axes, which, while comprehensive, omit higher-order reasoning over goals, causality, and multi-agent dynamics.
-Our experiments focus on short to medium context lengths; scaling to long-horizon rollouts remains an open challenge, especially under limited supervision.
-Although compute-efficient, training could be further improved with adaptive curricula or progressive tuning.
-Finally, like all pretrained VLMs, universe may reflect dataset biases and underperform on rare or ambiguous behaviors.
-Supplementary Materials
-[…]
-[…]
+Our large-scale study demonstrates that universe matches the performance of all task-specific checkpoints using a single unified model and aligns closely with human judgments, establishing it as a scalable, semantics-aware evaluator for evaluating world models, particularly when explicit ground truth is unavailable or costly to obtain.
+Limitations.
+While our experiments focus on simulated environments, chosen both for their ground-truth availability and their direct relevance to large video-game and interactive-entertainment industries Kanervisto et al.
+(2025), evaluating universe in real-world settings remains an important next step.
+Our protocol focuses on foundational semantic tasks, and extending it to cover higher-level reasoning represents an important direction for future work.
+Scaling universe to long-horizon rollouts is also challenging, as fine-grained reasoning becomes harder with larger visual context.
+Our frame-sampling analysis suggests that more intelligent sampling or hierarchical summarization could address this, and we plan to explore such strategies.
+Finally, as with all pretrained VLMs, universe may inherit dataset biases and exhibit reduced reliability on ambiguous cases.
+While we examined them partially during evaluation, a deeper investigation into bias propagation is an exciting research direction for future work.
+Appendix
 Appendix A Broader Impact
-[…]
-This capability has direct implications for high-impact domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77], where world models simulate environment dynamics and support downstream control and generalization.
+As world models become integral to simulation, planning, and decision-making in interactive environments, evaluation remains a key bottleneck for both research progress and safe deployment.
+We address this challenge by introducing a unified, sample-efficient framework for evaluating world model rollouts using adapted VLMs, designed for fine-grained, temporally grounded, and semantically coherent assessment.
+This capability has direct implications for high-impact domains such as neural game engines (Kanervisto et al., 2025; Guo et al., 2025; Gao et al., 2025; Chen et al., 2025), embodied AI (Du et al., ; Yang et al., 2024), and autonomous driving (Russell et al., 2025; Hu et al., 2023a; Ni et al., 2025), where world models simulate environment dynamics and support downstream control and generalization.
 In such contexts, precise and interpretable evaluation is critical not only for benchmarking, but also for diagnosing failure modes and ensuring alignment with intended behaviors.
-[…]
+By reducing dependence on human annotation and task-specific fine-tuning, universe offers a scalable alternative that lowers the computational and environmental costs of rollout evaluation.
+However, reliance on automated evaluators introduces risks: adapted VLMs may inherit biases from pretraining, struggle under distributional shift, or yield unreliable judgments in edge cases.
+These risks are amplified in safety-critical settings, where miscalibrated evaluations can propagate downstream errors.
 We therefore advocate for cautious deployment, accompanied by human oversight, rigorous validation, and transparent reporting.
@@
 To support reproducibility and facilitate future research, we provide detailed instructions for reproducing all main experiments.
-Detailed descriptions of model architectures, training procedures, and dataset construction are provided in Section 4 and Appendix D.
+Detailed descriptions of model architectures, training procedures, and dataset construction are provided in Section 4 and Appendix E.
+A high-level overview of the overall implementation framework is included in Appendix C.1.
 All experiments have been repeated for three runs.
 Plots and tables with quantitative results show the standard deviation across these runs.
-[…]
+Use of Existing Assets.
+We experiment with a range of open-weight VLMs, including three PaliGemma variants (version 1 (3B) (Beyer et al., 2024) and version 2 (3B and 10B) (Steiner et al., 2024)), VideoLLaMA3 (2B, 7B) (Boqiang Zhang, 2025), and CLIP (Radford et al., 2021) with the following vision encoder configurations: ViT-B/32, ViT-B/16, ViT-L/14, and ViT-L/14 with $336\times 336$ resolution.
+universe is built on top of PaliGemma v2 (3B), using publicly released checkpoints for initialization.
+Further architectural and implementation details are provided in Appendix E.1.
+For our software stack, we use Matplotlib (Hunter, 2007) for plotting, NumPy (Harris et al., 2020) for data handling, openCV (Bradski, 2000), FFmpeg (Tomar, 2006) and PIL (Umesh, 2012) for video and image processing, and NLTK (Bird & Loper, 2004) for text processing.
+Parameter-efficient fine-tuning is implemented using the PEFT library (Mangrulkar et al., 2022).
+We log our experiments using Weights and Biases (Biewald, 2020).
+Use of Large Language Models.
+Portions of the manuscript were polished with the assistance of large language models (LLMs).
+The use of LLMs was limited to only improving readability and style; all ideas, and experimental designs were developed by the authors.
 Compute Resources.
@@
 In total, all experiments amounted to approximately 5,153 GPU-days, equivalent to 14.12 GPU-years.
-Appendix C Dataset
-[…]
-C.1 Construction Process
+Appendix C universe: Additional Details
+C.1 Implementation Overview
+This section outlines the implementation of universe in Python, presented as high-level pseudocode.
+The system is structured around two main stages: (i) Adaptation:fine-tuning a VLM on task-specific question-answer (QA) supervision derived from ground truth; (ii) Evaluation:using the adapted model to assess new rollouts via structured, prompt-based recognition tasks.
+Adaptation Pipeline.
+The adaptation stage can be implemented as two modules: AdaptationDatasetBuilder and VLMAdapter.
+AdaptationDatasetBuilder.
+This class constructs an adaptation dataset from raw ground truth data, initialized via load_ground_truth_data (see Section 3 and Appendix D).
+The core method, build, takes four arguments: alpha_task, which specifies the task mixture ratio; beta_format, which controls the distribution over QA prompt formats; context_length, which determines the number of frames per QA instance; and sampling_strategy, which defines how frames are sampled from rollouts.
+The builder first applies stratified_sample to select a subset of annotated samples that match the specified configuration.
+For each sample, it invokes _sample_visual_context to extract the relevant frames, and constructs a triplet consisting of frames, question, and answer.
+VLMAdapter.
+This class applies an adaptation strategy to a base VLM, passed via the base_vlm argument.
+Given an adaptation dataset adaptation_data, a tuning strategy specified by the strategy parameter, and a fixed number of training steps num_steps, the adapter trains the model by iteratively sampling a batch, computing the loss via compute_loss, and applying updates with update_model.
+Evaluation Pipeline.
+The evaluation stage can be implemented via two additional modules: RolloutsGenerator and Universe.
+RolloutsGenerator.
+This component autoregressively samples rollout trajectories from a world model (textttworld_model).
+Given an initial observation o_initial and an action sequence a_seq, the rollout method generates a sequence of predicted observations by maintaining lists of past observations (o_lt) and actions (a_lt).
+At each timestep, it calls predict_next_observation to obtain the next predicted frame, appends it to the rollout sequence o_seq, and continues until timestamps is reached.
+This process produces a full trajectory simulating environment dynamics.
+Universe.
+This module serves as the inference engine of our framework.
+It wraps an adapted VLM passed via adapted_vlm.
+Given a generated rollout and an evaluation specification, the method evaluate_rollout constructs a prompt using generate_question, parameterized by a recognition target and complexity level.
+It then calls evaluate, which queries the VLM with the resulting rollout and question, returning the model’s answer.
+C.2 Adaptation to New Domains
+To adapt universe to a new environment, one could collect a small set of reference trajectories that provide ground-truth observations and actions for the evaluation dimensions of interest.
+Question and answer templates can then be instantiated from these ground-truth signals to construct a mixed-supervision adaptation dataset.
+In practice, two complementary routes exist for obtaining such reference data: (i) partnerships with game studios (as in the present work) provide high-quality logs and metadata but are resource-intensive, (ii) bootstrapping from publicly available gameplay video datasets, where generative models or synthetic data can supply initial labels.
+At inference time, the evaluator could process rollouts by emitting natural-language responses along the specified dimensions.
+These outputs may be used either as structured feedback for improving the world model or mapped to numerical scores and aggregated into a single trajectory-level quality measure.
+Although our experiments use 14-frame inputs, this is a design choice rather than a conceptual constraint: the same procedure can extend to longer rollouts by increasing the visual context window or applying a sliding-window scheme over 14-frame segments.
+Overall, deploying universe as a standalone evaluator would require: (i) a small adaptation dataset of reference trajectories with ground-truth labels, (ii) the task specification, and (iii) the lightweight fine-tuning recipe.
+Once adapted, the evaluator could provide dimension-wise assessments that can be aggregated or used as feedback in downstream world-model training.
+Appendix D Dataset
+This section details the construction and release of the dataset used to adapt VLMs for fine-grained evaluation of world model rollouts.
+We curate a realistic, human-centered dataset derived from actual gameplay in a complex multi-agent environment.
+Designed to provide temporally grounded and semantically structured supervision, the dataset aligns with the downstream evaluation setting and supports adaptation to both action and character recognition tasks across all QA formats.
+We describe the data construction pipeline, QA generation process, and release format below.
+D.1 Construction Process
 The ground truth dataset for adapting the evaluator (see Section 3) was developed in collaboration with Ninja Theory using human gameplay recordings from Bleeding Edge, a 4v4 multiplayer combat game.
@@
 Next, for each sequence of frames $f^{(1:L)}\in\mathcal{V}$, we use the associated control log $c^{(1:L)}$ to extract action information and the metadata $m$ to obtain character-related attributes.
-These are combined to generate a structured natural language description via $d=\texttt{Describe}(c^{(1:L)},m)$.
+These are combined to generate a structured natural language description via $d=\texttt{Describe}(c^{(1:L)},m)$, where Describe is a rule-based procedure that transforms the logged actions and metadata into textual descriptions used for constructing the QA supervision.
 This yields a set of paired video–text examples: $\mathcal{Z}=\left\{(f^{(1:L)},d)\mid f^{(1:L)}\in\mathcal{V}\right\}$.
   3. (iii)
-    Question-Answer Pair Construction.
-Finally, we generate six QA pairs per clip, spanning two predefined tasks (AR and CR), each instantiated in three question formats: binary, multiple-choice, and open-ended.
-To enable this, we define task-specific answer spaces using $\texttt{GetAnswerSpace}(\mathcal{Z})$, which returns $\mathcal{Y}_{\text{AR}}$ for action categories and $\mathcal{Y}_{\text{CR}}$ for character identities, based on all video–text pairs in $\mathcal{Z}$.
-For each clip, we extract the task-specific ground-truth answer from the corresponding description as $y=\texttt{ExtractLabel}(d,t)$, where $t\in\{\text{AR},\text{CR}\}$.
-Each QA format is constructed as follows: (i) Binary: Two binary question-answer pairs are generated per instance using FormatBinaryPrompt.
-The positive question $Q^{\text{pos}}$ is constructed using the correct label $y\in\mathcal{Y}^{(t)}$ and paired with the positive answer $A^{\text{pos}}$.
-The negative question $Q^{\text{neg}}$ is constructed using an incorrect label $\tilde{y}\sim\texttt{SampleDistractor}(\mathcal{Y}^{(t)}\setminus\{y\})$ and paired with the negative answer $A^{\text{neg}}$.
-(ii) Multiple-Choice: A question $Q$ is generated using the full set of candidate options, formatted via FormatOptions$(\mathcal{Y}_{t})$.
-The question is constructed with FormatMCPrompt$(t,O)$ and paired with the correct answer $y\in\mathcal{Y}_{t}$.
-(iii) Open-Ended: A free-form question $Q$ is generated using FormatOEPrompt$(t)$, prompting the model to produce the correct label $y\in\mathcal{Y}_{t}$ without access to predefined answer choices.
+[…]
 The final dataset is represented as $\mathcal{D}=\{(f_{i}^{(1:L)},QA_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $f^{(1:L)}$ is a video clip and $QA=\{(Q_{j},A_{j})\}_{j=1}^{6}$ is the associated set of question–answer pairs, covering all combinations of three question formats (binary, multiple-choice, open-ended) and two tasks (Action Recognition and Character Recognition).
@@
  $Q^{pos}\leftarrow\texttt{FormatBinaryPrompt}(t,y)$
-      $Q^{neg}\leftarrow\texttt{FormatBinaryPrompt}(t,\tilde{y})$
+[…]
  return $\{(Q^{pos},A^{pos}),\;(Q_{\tilde{y}},A^{neg})\}$
@@
 Algorithm 1 Dataset Construction Process
-Appendix D Experimental Details
+Appendix E Experimental Details
 In this section, we provide a detailed description of the dataset preparation process, model architecture, prompt templates, training procedure.
 Additionally, we provide an overview of all results presented in the main paper in numerical table form, an report additional experimental results leveraging alternate fine-tuning solutions.
-D.1 Model
+E.1 Model
 This section provides extended details on the architecture, pretraining configuration, and input formatting of the vision-language models used in our experiments.
-Our primary backbone is PaliGemma [10, 91].
+Our primary backbone is PaliGemma (Beyer et al., 2024; Steiner et al., 2024).
 Table 1: Detailed architecture of the PaliGemma model, comprising a SigLIP-So400m vision tower, a multimodal projection head, and a Gemma-based language decoder.
@@
 LM Head Linear(2304 $\rightarrow$ 257216, bias=False)
-D.1.1 Overview
-[…]
+E.1.1 Overview
+PaliGemma is a VLM that processes both images and text as input and autoregressively generates natural language output.
+It follows the training paradigm of PaLI-3 (Chen et al., 2023), combining a ViT-based vision encoder (Dosovitskiy et al., 2021) with a decoder-only Transformer language model.
+The model is publicly available (Wolf et al., 2019).
+The architecture is fully modular, comprising three parameterized components: (i) Vision encoder($\mathcal{M}_{V}$): based on SigLIP (Zhai et al., 2023), specifically the “shape optimized” So400m (Alabdulmohsin et al., 2023).
+(ii) Multimodal projection head($\mathcal{M}_{P}$): a single linear layer for projecting visual features into the language decoder’s embedding space.
+(iii) Language decoder($\mathcal{M}_{L}$): a Transformer-based autoregressive model from the Gemma family (Mesnard et al., 2024; Rivière et al., 2024).
+Below, we discuss the architecture in more details, the general layer-level overview is also provided in Table 1.
 Vision Encoder: SigLIP-So400m.
@@
 The language module $\mathcal{M}_{L}$ is a decoder-only Transformer with 26 layers and 2304-dimensional hidden states.
-Token embeddings are learned over a vocabulary of 257,216 tokens, encoded using the SentencePiece tokenizer [58].
+Token embeddings are learned over a vocabulary of 257,216 tokens, encoded using the SentencePiece tokenizer (Kudo & Richardson, 2018).
 Each Transformer block contains a self-attention mechanism with separate linear projections for queries, keys, and values.
-The MLP block follows a gated architecture, where the input is processed through parallel down projection and gating projection layers, modulated by a GELU-Tanh activation [44], combined via elementwise multiplication, and then passed through an up projection to return to the model’s hidden dimension.
+The MLP block follows a gated architecture, where the input is processed through parallel down projection and gating projection layers, modulated by a GELU-Tanh activation (Hendrycks & Gimpel, 2016), combined via elementwise multiplication, and then passed through an up projection to return to the model’s hidden dimension.
 RMSNorm is applied before and after both attention and MLP sublayers to stabilize training.
@@
 Output tokens are produced via a tied language modeling head that projects back to the vocabulary space.
-D.1.2 Configurations
+E.1.2 Configurations
 Table 2 summarizes the architecture components and parameter counts of the PaliGemma configurations available for experimentation.
@@
 PG 3 Gemma 2 9B, pre-trained on 8T tokens 9.7B
-D.1.3 Prompt Format
+E.1.3 Prompt Format
 To generate textual responses, we adopt a unified prompt format for the decoder.
@@
 The overall input schema is as follows:
-$S=\;$ $\underbrace{\texttt{<image>}_{1}^{(1)},\ldots,\texttt{<image>}_{p}^{(1)},\;\ldots,\;\texttt{<image>}_{1}^{(n)},\ldots,\texttt{<image>}_{p}^{(n)}}_{S_{\mathcal{I}}\text{: Visual tokens from }n\text{ frames, each represented as }p\text{ patches}}$
+[…]
 $\underbrace{\texttt{<BOS>},\;\texttt{answer en},\;\texttt{<QUESTION>},\;\texttt{<SEP>}}_{S_{\mathcal{T}}^{\text{PREF}}\text{: Prefix (cue + question)}}$
@@
 The suffix $S_{\mathcal{T}}^{\text{SUFF}}$ contains the target answer, terminated with <EOS> and padded with <PAD> tokens for batching.
-D.1.4 Pretraining Data and Filtering
-PaliGemma is pretrained on a mixture of large-scale vision-language datasets, including WebLI [23], CC3M-35L [87], VQ2A-CC3M-35L [19], OpenImages [80], and WIT [90].
+E.1.4 Pretraining Data and Filtering
+PaliGemma is pretrained on a mixture of large-scale vision-language datasets, including WebLI (Chen et al., 2022b), CC3M-35L (Sharma et al., 2018), VQ2A-CC3M-35L (Changpinyo et al., 2022), OpenImages (Piergiovanni et al., 2022), and WIT (Srinivasan et al., 2021).
 Data quality and safety are maintained through pornographic content filtering, text safety and toxicity filtering, and privacy-preserving measures.
-D.2 Evaluation Metrics
+E.2 Evaluation Metrics
 In this section, we provide additional details on metrics used for quantitative evaluation.
@@
 For binary questions, we compute the metric on the bigram level, while for multiple-choice and open-ended formats, we use trigram-level evaluation.
-Formally, let $G$ and $R$ denote the sets of $n$-grams in the generated and reference answers, respectively. Precision and recall are defined as:
-$P=\frac{|G\cap R|}{|G|},\quad R=\frac{|G\cap R|}{|R|}$ (3)
-where $|G\cap R|$ counts overlapping $n$-grams. The ROUGE score is then computed as:
+Formally, let $G$ and $GT$ denote the sets of $n$-grams in the generated and ground truth answers, respectively. Precision and recall are defined as:
+$P=\frac{|G\cap GT|}{|G|},\quad R=\frac{|G\cap GT|}{|GT|}$ (3)
+where $|G\cap GT|$ counts overlapping $n$-grams. The ROUGE score is then computed as:
 $\text{ROUGE}=2\times\frac{P\times R}{P+R}$ (4)
 Together, these metrics provide a robust view of model performance: EM reflects exact correctness, while ROUGE provides a softer measure of semantic fidelity, particularly useful for evaluating open-ended generations.
-D.3 Results
+E.3 Results
 Table 3: Summary of hyperparameters used in our experiments.
@@
 Gradient accumulation steps 4
-Optimizer AdamW [67]
+Optimizer AdamW Loshchilov & Hutter (2019)
 Learning rate $5\times 10^{-5}$, cosine annealing
@@
 Gradient clipping Global norm, threshold 1.0
-VLM backbone PaliGemma 2 (3B) [10]
+VLM backbone PaliGemma 2 (3B) Beyer et al. (2024)
 Hyperparameters.
@@
 Models were trained for 1–10 epochs depending on task and setting.
-Optimization is performed using AdamW [67] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, a base learning rate of $5\times 10^{-5}$, and weight decay of $1\times 10^{-6}$.
-We use cosine learning rate annealing [68] with a linear warmup over the first 10% of training steps.
+Optimization is performed using AdamW (Loshchilov & Hutter, 2019) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, a base learning rate of $5\times 10^{-5}$, and weight decay of $1\times 10^{-6}$.
+We use cosine learning rate annealing (Loshchilov & Hutter, 2022) with a linear warmup over the first 10% of training steps.
 To stabilize training, we apply gradient clipping with a global norm threshold of 1.0.
-All models use PaliGemma 2 (3B) [10] as the vision-language backbone unless otherwise noted.
+All models use PaliGemma 2 (3B) (Beyer et al., 2024) as the vision-language backbone unless otherwise noted.
 We vary the number of input frames between 1 and 8 depending on task, and all images are resized to a fixed resolution of $224\times 224$.
@@
 Fr Model Binary MC OE Binary MC OE
-1 PG 1 3B 50.43 $\pm$ 0.13 08.12 $\pm$ 0.02 10.83 $\pm$ 0.01 50.73 $\pm$ 0.38 00.46 $\pm$ 0.06 $0$$\pm$ 0.00
+1 PG 1 3B 50.43 $\pm$ 0.13 08.12 $\pm$ 0.02 10.83 $\pm$ 0.01 50.73 $\pm$ 0.38 00.46 $\pm$ 0.06 00.00 $\pm$ 0.00
 PG 2 3B 44.69 $\pm$ 0.03 09.30 $\pm$ 0.17 12.64 $\pm$ 0.01 48.58 $\pm$ 0.07 00.28 $\pm$ 0.06 00.01 $\pm$ 0.00
-PG 2 10B 50.04 $\pm$ 0.03 26.98 $\pm$ 0.00 12.35 $\pm$ 0.21 50.08 $\pm$ 0.07 08.33 $\pm$ 0.50 $0$$\pm$ 0.00
+PG 2 10B 50.04 $\pm$ 0.03 26.98 $\pm$ 0.00 12.35 $\pm$ 0.21 50.08 $\pm$ 0.07 08.33 $\pm$ 0.50 00.00 $\pm$ 0.00
 VL3-2B 03.24 $\pm$ 0.00 18.52 $\pm$ 0.06 06.27 $\pm$ 0.05 08.76 $\pm$ 0.04 03.44 $\pm$ 0.08 00.50 $\pm$ 0.01
 VL3-7B 45.02 $\pm$ 0.28 15.53 $\pm$ 0.05 06.54 $\pm$ 0.04 39.09 $\pm$ 0.73 06.21 $\pm$ 0.05 00.51 $\pm$ 0.02
-8 PG 1 3B 51.67 $\pm$ 0.02 10.68 $\pm$ 0.00 10.32 $\pm$ 0.00 51.39 $\pm$ 0.07 00.25 $\pm$ 0.00 $0$$\pm$ 0.00
+8 PG 1 3B 51.67 $\pm$ 0.02 10.68 $\pm$ 0.00 10.32 $\pm$ 0.00 51.39 $\pm$ 0.07 00.25 $\pm$ 0.00 00.00 $\pm$ 0.00
 PG 2 3B 47.61 $\pm$ 0.19 06.73 $\pm$ 0.04 14.52 $\pm$ 0.00 48.37 $\pm$ 0.18 00.03 $\pm$ 0.00 00.01 $\pm$ 0.00
-PG 2 10B 50.02 $\pm$ 0.06 26.93 $\pm$ 0.01 12.12 $\pm$ 0.00 50.09 $\pm$ 0.06 00.22 $\pm$ 0.00 $0$$\pm$ 0.00
+PG 2 10B 50.02 $\pm$ 0.06 26.93 $\pm$ 0.01 12.12 $\pm$ 0.00 50.09 $\pm$ 0.06 00.22 $\pm$ 0.00 00.00 $\pm$ 0.00
 VL3-2B 13.92 $\pm$ 0.13 03.47 $\pm$ 0.02 00.32 $\pm$ 0.01 13.92 $\pm$ 0.13 03.46 $\pm$ 0.04 00.32 $\pm$ 0.01
@@
 $\mathcal{F}_{all}$ 74.35 $\pm$ 1.37 74.35 $\pm$ 1.37 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00 13.13 $\pm$ 0.00 27.57 $\pm$ 0.00
-$\mathcal{F}_{LoRA}$ 44.66 $\pm$ 0.21 44.66 $\pm$ 0.21 00.02 $\pm$ 0.01 09.21 $\pm$ 0.01 $0$$\pm$ 0.00 12.49 $\pm$ 0.00
+$\mathcal{F}_{LoRA}$ 44.66 $\pm$ 0.21 44.66 $\pm$ 0.21 00.02 $\pm$ 0.01 09.21 $\pm$ 0.01 00.00 $\pm$ 0.00 12.49 $\pm$ 0.00
 Character Recognition
@@
 $\mathcal{F}_{all}$ 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 97.67 $\pm$ 0.06 97.67 $\pm$ 0.06 96.55 $\pm$ 0.01 96.55 $\pm$ 0.01
-$\mathcal{F}_{LoRA}$ 48.76 $\pm$ 0.00 48.76 $\pm$ 0.00 $0$$\pm$ 0.00 00.32 $\pm$ 0.00 $0$$\pm$ 0.00 00.01 $\pm$ 0.01
+$\mathcal{F}_{LoRA}$ 48.76 $\pm$ 0.00 48.76 $\pm$ 0.00 00.00 $\pm$ 0.00 00.32 $\pm$ 0.00 00.00 $\pm$ 0.00 00.01 $\pm$ 0.01
 Table 6: Supervision and Temporal Context: Training budget analysis for Character Recognition.
@@
 Fr EM ROUGE EM ROUGE EM ROUGE
-First-N 1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
+First-N
+1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
 2 84.42 $\pm$ 0.06 84.42 $\pm$ 0.06 65.53 $\pm$ 0.27 72.03 $\pm$ 0.06 65.38 $\pm$ 0.06 71.74 $\pm$ 0.23
@@
 8 93.85 $\pm$ 0.28 93.85 $\pm$ 0.28 88.89 $\pm$ 0.14 93.40 $\pm$ 0.76 87.80 $\pm$ 0.20 92.23 $\pm$ 0.16
-Uniform-N 1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
+Uniform-N
+1 83.97 $\pm$ 0.02 83.97 $\pm$ 0.02 61.43 $\pm$ 0.58 68.05 $\pm$ 0.70 61.68 $\pm$ 0.35 68.46 $\pm$ 0.19
 2 90.47 $\pm$ 0.62 90.47 $\pm$ 0.62 83.93 $\pm$ 0.04 88.36 $\pm$ 0.08 82.68 $\pm$ 0.19 87.33 $\pm$ 0.01
@@
 0.05 0.15 0.80 71.79 $\pm$ 0.01 71.79 $\pm$ 0.01 98.22 $\pm$ 0.00 98.22 $\pm$ 0.00 98.57 $\pm$ 0.00 98.57 $\pm$ 0.00
-Appendix E Human Annotation Study
+Appendix F Human Annotation Study
 This section provides full details of our human annotation study, including rollout generation, annotation procedures, inter-annotator agreement, and evaluation metrics.
 The goal is to validate the adapted VLM’s fine-grained predictions on generated video rollouts.
-E.1 Study Design
+F.1 Study Design
 Table 13: Annotation instructions provided to human raters as part of the study.
@@
 Selected World Models.
-For our study, we select two autoregressive world models of different scales, both based on the WHAM architecture [56], a publicly available world model.
-These models are trained to model sequences of visual frames and controller actions, without any textual supervision.
-Each world model is a decoder-only transformer [81, 98] trained to autoregressively predict discrete tokens representing visual observations and actions.
-Visual frames are first encoded using a VQGAN [32], while joystick actions are tokenized using a learned discretization scheme based on action bucketization [55].
-The model is trained to predict the next token in the sequence, conditioned on prior visual and action tokens.
-Specifically, we focus on two versions of WHAM with differing model capacities and training environments:
-  • WHAM 140M: A 140M-parameter model trained for 100K steps on gameplay from a single environment (Environment A / Skygarden) at $128\times 128$ resolution.
-  • WHAM 1.6B: A 1.6B-parameter model trained for 200K steps on gameplay from seven environments (Environments A–G, including Skygarden) at $300\times 180$ resolution.
+For our study, we aim to evaluate rollouts generated across different model scales, training diversities, and output resolutions, while keeping the underlying architecture general enough to apply broadly.
+To this end, we select two autoregressive world models (Kanervisto et al., 2025).
+The autoregressive formulation offers a flexible and widely adopted framework, and is the basis for many state-of-the-art private world models.
+The selected models generate sequences of visual frames and controller actions without textual supervision, using a decoder-only transformer (Radford et al., 2019; Vaswani et al., 2017) trained autoregressively on discrete tokens.
+Visual frames are encoded with a VQGAN (Esser et al., 2021), while joystick actions are tokenized using a learned discretization scheme based on action bucketization (Kanervisto et al., 2020).
+Figure 11: Annotation interface example.
+Each instance includes a video clip, task instructions, and a table with: Question (generated via evaluation protocol), Response (VLM output), and Score (human-assigned label).
+Each world model is implemented as a decoder-only transformer (Radford et al., 2019; Vaswani et al., 2017), trained to predict discrete tokens representing visual observations and actions.
+Visual frames are first compressed with a VQGAN (Esser et al., 2021), while joystick actions are tokenized using a learned discretization scheme based on action bucketization (Kanervisto et al., 2020).
+The objective is next-token prediction conditioned on prior visual and action tokens.
+We focus on the following model variants that differ in capacity, training diversity, and output resolution: (i) a large-scale model: 1.6B parameters, trained for 200K steps on gameplay from seven diverse environments (including Skygarden) at $300\times 180$, and (ii) a smaller model: 140M parameters, trained for 100K steps on gameplay from a single environment (Skygarden) at $128\times 128$ resolution.
 Rollouts Generation.
 Rollout generation follows a consistent protocol for both world models: at inference time, the model is conditioned on 1 second of ground-truth gameplay (visual and action tokens), after which it generates 10 seconds of future gameplay conditioned only on a sequence of held-out controller actions.
+For rollout generation, we use reference trajectories from the game dataset, which include time-aligned controller inputs and image frames.
+For each generated rollout, we randomly sample a 1-second context window (video frames and corresponding action tokens) from these trajectories.
+The world model is then conditioned on this context, which consists of an interleaved sequence of tokenized images and actions, and proceeds autoregressively to generate latent image tokens and discretized action tokens.
 The generated rollout is then split into 14-frame chunks.
-This setup enables a comprehensive analysis of the universe’s evaluation capabilities across two axes: (i) in-domain performance: evaluating on Skygarden (Environment A), the environment used for fine-tuning; (ii) generalization: assessing performance on six unseen environments (Environments B–G).
+This setup enables a comprehensive analysis of the universe’s evaluation capabilities across two axes: (i) in-domain performance: evaluating on Skygarden, the environment used for fine-tuning; (ii) generalization: assessing performance on six unseen environments.
 It also allows comparison across generation quality and model capacity.
 We generate 82 rollouts for each model-environment setting, resulting in 656 rollouts in total.
-Figure 10: Annotation interface example.
-Each instance includes a video clip, task instructions, and a table with: Question (generated via evaluation protocol), Response (VLM output), and Score (human-assigned label).
+Figure 12: Randomly sampled examples of rollouts excluded from the human evaluation study.
 Rollout Filtering.
-To ensure quality and clarity, we filtered out rollouts that: (i) had no visible agents, (ii) featured stationary agents, (iii) were taken from early uninformative environment segments, or (iv) had significant visual obstruction.
-We also excluded sequences containing more than four characters to reduce annotation ambiguity.
-universe Response Generation.
+To keep evaluation focused on challenging sequences, we remove only those rollouts that are either uninterpretable or uninformative.
+Specifically, annotators discard: (i) early segments with no meaningful interactions; (ii) rollouts dominated by occlusion; and (iii) sequences with no visible agents.
+These criteria target cases where neither humans nor models can extract actionable dynamics, while retaining all visually imperfect, ambiguous, and failure-prone generations.
+Thus, filtering eliminates only sequences that provide no evaluable signal.
+Figure 12 shows randomly sampled examples from each excluded category.
+Figure 13: Randomly sampled examples of rollouts retained for human evaluation.
+Complexity of Retained Rollouts.
+Despite filtering out uninterpretable/uninformative sequences, many retained rollouts exhibit visual imperfections that make evaluation challenging.
+Annotators highlighted recurring generation failures including: (i) skin and appearance drift, where model-generated agent skins diverge from canonical silhouettes; (ii) missing or corrupted assets, such as hoverboards implied only through motion cues; and (iii) HUD-induced occlusion, where oversized or misplaced UI elements block important scene content.
+These imperfections arise from the underlying world models and introduce ambiguity while still permitting meaningful assessment.
+Figure 13 illustrates representative failure-prone cases, emphasizing that universe is evaluated on realistic, imperfect rollouts rather than sanitized inputs.
+universe’s Response Generation.
 To obtain responses from universe, we provide it with a video segment (resized to match the evaluator’s input resolution) along with its corresponding question.
@@
 In cases where all five responses are unique (i.e., no majority), one response is selected at random.
-The resulting dataset comprises rollouts from 8 model–environment pairs: rollouts generated by WHAM 140M on Environment A (Skygarden), and rollouts generated by WHAM 1.6B across seven distinct environments (Environments A–G).
+The resulting dataset comprises rollouts from 8 settings: rollouts generated by a smaller model on Skygarden, and rollouts generated by a larger model across seven distinct environments.
 For each model–environment pair, we sample 30 rollouts.
@@
 Each of the resulting 1,440 QA instances was rated by 3 annotators, yielding 4,320 total human judgments.
-E.2 Evaluation Metrics
+F.2 Evaluation Metrics
 We report two accuracy-based metrics using the adjudicated labels:
@@
 The total number of valid examples excludes QA pairs marked as Unclear by at least one annotator.
-Model Env. Valid QA Pairs Cohen’s $\kappa$
-WHAM 140M A 24 0.79
-WHAM 1.6B A 29 0.91
-B 28 0.67
-C 28 0.74
-D 29 0.87
-E 30 0.67
-F 29 0.61
-G 30 0.59
+Setting Valid QA Pairs Cohen’s $\kappa$
+1 29 0.91
+2 28 0.67
+3 28 0.74
+4 29 0.87
+5 30 0.67
+6 29 0.61
+7 30 0.59
+8 24 0.79
 Sample Size Justification.
 We annotate 30 rollouts per model–environment pair.
-Assuming a standard deviation of $\sigma\approx 0.2$ and a 95% confidence level, the confidence interval (CI) width is given by CI $\text{CI Width}=z_{\frac{1-C}{2}}\cdot\frac{\sigma}{\sqrt{n}}$.
-This yields an estimated CI of $\sim$7.1for individual model–environment pairs ($n=30$), and $\sim$2.5when aggregating across all eight pairs ($n=240$), offering sufficient precision for comparative evaluation.
+Assuming a standard deviation of $\sigma\approx 0.2$ and a 95% confidence level, the confidence interval (CI) width is given by $\text{CI Width}=z_{\frac{1-C}{2}}\cdot\frac{\sigma}{\sqrt{n}}$.
+This yields an estimated CI of $\sim$7.1% for individual model–environment pairs ($n=30$), and $\sim$2.5% when aggregating across all eight pairs ($n=240$), offering sufficient precision for comparative evaluation.
 Table 15: Graded/strict accuracy of universe on Action and Character Recognition tasks, evaluated by human annotators across different environments and question formats.
@@
 Action Recognition Character Recognition
-Model Env. Binary MC OE Binary MC OE
-WHAM 140M A 92.9 / 89.3 35.7 / 32.1 41.1 / 39.3 85.7 / 85.7 10.7 / 10.7 60.7 / 60.7
-WHAM 1.6B A 98.3 / 96.7 51.7 / 46.7 75.0 / 73.3 93.3 / 93.3 83.3 / 83.3 93.3 / 93.3
-B 96.7 / 96.7 60.0 / 60.0 65.0 / 60.0 99.9 / 99.9 90.0 / 90.0 93.3 / 93.3
-C 96.7 / 96.7 63.3 / 63.3 80.0 / 80.0 99.9 / 99.9 86.7 / 86.7 93.3 / 93.3
-D 93.3 / 93.3 43.3 / 43.3 73.3 / 73.3 96.7 / 96.7 96.7 / 96.7 99.9 / 99.9
-E 80.0 / 76.7 76.7 / 73.3 93.3 / 93.3 96.7 / 96.7 99.9 / 99.9 99.8 / 99.8
-F 71.7 / 70.0 56.7 / 56.7 75.0 / 70.0 96.7 / 96.7 93.3 / 93.3 96.7 / 96.7
-G 68.3 / 66.7 50.0 / 46.7 80.0 / 76.7 93.3 / 93.3 90.0 / 90.0 96.7 / 96.7
-E.3 Results
-[…]
+Setting Binary MC OE Binary MC OE
+1 98.3 / 96.7 51.7 / 46.7 75.0 / 73.3 93.3 / 93.3 83.3 / 83.3 93.3 / 93.3
+2 96.7 / 96.7 60.0 / 60.0 65.0 / 60.0 99.9 / 99.9 90.0 / 90.0 93.3 / 93.3
+3 96.7 / 96.7 63.3 / 63.3 80.0 / 80.0 99.9 / 99.9 86.7 / 86.7 93.3 / 93.3
+4 93.3 / 93.3 43.3 / 43.3 73.3 / 73.3 96.7 / 96.7 96.7 / 96.7 99.9 / 99.9
+5 80.0 / 76.7 76.7 / 73.3 93.3 / 93.3 96.7 / 96.7 99.9 / 99.9 99.8 / 99.8
+6 71.7 / 70.0 56.7 / 56.7 75.0 / 70.0 96.7 / 96.7 93.3 / 93.3 96.7 / 96.7
+7 68.3 / 66.7 50.0 / 46.7 80.0 / 76.7 93.3 / 93.3 90.0 / 90.0 96.7 / 96.7
+8 92.9 / 89.3 35.7 / 32.1 41.1 / 39.3 85.7 / 85.7 10.7 / 10.7 60.7 / 60.7
+F.3 Results
+Table 15 reports graded and strict accuracy across environments, recognition targets (Action and Character Recognition), and question formats (Binary, Multiple-Choice, Open-Ended).
+We observe a clear gap in performance between rollouts generated by the two world models.
+universe struggles with outputs from WHAM 140M, achieving substantially lower accuracy compared to WHAM 1.6B.
+This is likely due to a mismatch in image resolution: WHAM 140M generates frames at $128\times 128$ resolution, which must be upsampled to the universe’s expected input of $224\times 224$.
+Despite resizing, the resulting frames often lack sharpness, making actions and characters harder to recognize.
+In contrast, universe performs well on rollouts from WHAM 1.6B, even across diverse environments.
+On the in-domain setting (Environment A), the model achieves strong results—averaging 75.02% graded accuracy for AR and 90.00% for CR.
+When evaluating on the six unseen environments (Environments B–G), performance for AR drops slightly (from 75.02% to 73.52%), while CR remains stable or improves, suggesting strong generalization in character grounding and visual consistency tracking.
 Qualitative Examples. Figure 14 illustrates the diversity of generated rollouts across environments. WHAM 1.6B captures greater visual variation and scene composition compared to WHAM 140M.
-Figure 11: Representative frames from rollouts generated across seven environments.
-WHAM 140M (top row) was trained only on Skygarden (Environment A); WHAM 1.6B (rows 2-8) generalizes across seven environments.
-Appendix F Supplementary Experimental Results
+Figure 14: Representative frames from rollouts across the eight evaluation settings, spanning different environments, scales, and resolutions.
+Appendix G Supplementary Experimental Results
 This section presents additional experimental results that support the main findings but are omitted from the main paper for clarity and space.
@@
 While these results are not central to the unified evaluation framework proposed in the main text, they provide valuable insight into model selection, adaptation efficiency, and the limitations of standard evaluation proxies in our setting.
-F.1 Zero-Shot Performance of PaliGemma Models
+G.1 GPT-5 Performance
+Figure 15: Performance of GPT-5 on six randomly sampled binary action recognition questions.
+Each panel shows the trial prompt, the model’s response, and the ground-truth label.
+Despite the apparent simplicity of the setup, GPT-5 fails on 5/6 trials, underscoring the difficulty of the task and the need for task-specific adaptation.
+To demonstrate the complexity of the tasks that comprise our protocol, we conducted an evaluation of GPT-5 on randomly selected examples.
+We deliberately chose the simplest evaluation regime (binary action recognition) to test whether the model can succeed without adaptation.
+Figure 15 demonstrates that out of six random samples, GPT-5 produced incorrect answers in five cases. This consistent failure highlights the difficulty of the task.
+G.2 Zero-Shot Performance of PaliGemma Models
 Figure 16: Zero-shot evaluation results for PaliGemma variants across tasks, prompt formats, and visual context sizes.
@@
 We therefore adopt PaliGemma 2 3b as the default model for all subsequent adaptation experiments, balancing performance, compute efficiency, and architectural recency.
-F.2 CLIPScore Comparisons
+G.3 CLIPScore Comparisons
 Table 16: Zero-shot accuracy-based evaluation of CLIP models and baseline methods on Action and Character Recognition tasks using 1 and 8 input frames.
@@
 As such, we exclude CLIP-based scores from our primary comparisons and instead focus on adapted, generative VLM-based evaluators.
-F.3 Low-Rank Adaptation Comparisons
+G.4 Low-Rank Adaptation Comparisons
 This section presents an extended analysis of low-rank adaptation (LoRA) as a parameter-efficient strategy for adapting vision-language models to our protocol.
@@
 Accordingly, all results reported in the main paper use $r=8$, while extended comparisons with higher ranks are presented here for completeness.
+Appendix H Comparison with Existing Evaluation Baselines
+To validate that universe captures evaluation dimensions beyond existing automated metrics, we analyze its correlation with two representative baselines: Fréchet Video Distance (FVD), a standard distributional metric for video generation quality, and VBench, a comprehensive multi-dimensional benchmark for text-to-video evaluation.
+These comparisons assess whether universe’s focus on action-conditioned semantic alignment provides complementary signal to metrics designed primarily for perceptual quality and temporal coherence.
+H.1 Comparison with FVD
+Table 18: Pearson correlation between FVD and universe across experimental settings. None of the correlations are statistically significant (all $p>0.05$).
+Setting Action Recognition Character Recognition
+1 0.09 0.03
+2 -0.07 -0.08
+3 -0.17 0.06
+4 -0.03 -0.10
+5 0.07 -0.25
+6 -0.01 0.21
+7 0.32 0.03
+8 -0.10 0.33
+FVD Unterthiner et al.
+(2018) is a widely used reference-free metric for evaluating generative video models, defined as a distance between feature distributions of real and generated videos.
+Although commonly treated as a default quantitative proxy for rollout quality in world-model evaluation, FVD is agnostic to task semantics and action alignment—the dimensions universe is specifically designed to capture.
+Across the eight environments, raw FVD values are tightly clustered (means between 3.23 and 3.27 with standard deviations below 0.05), reflecting limited dynamic range and strong dataset-dependent scale effects.
+As a result, raw scores are not directly comparable across settings, making correlations a more informative basis for analysis.
+We therefore assess the relationship between the two metrics by computing Pearson correlations between FVD and universe’s human-aligned scores across all environments for both Action Recognition and Character Recognition tasks.
+Results.
+For each environment (Settings 1–8) and each task, we obtain FVD scores on the same set of rollouts evaluated by universe, report the scores and compute the correlation between the two.
+The resulting coefficients are reported in Table 18.
+Correlations are small in magnitude and unstable in sign, ranging from 0.25-0.25 to $0.33$, with most values close to zero.
+The largest positive correlations are modest (Action Recognition, Setting 7: $r=0.32$; Character Recognition, Setting 8: $r=0.33$), and several settings exhibit weak negative correlations (e.g., Character Recognition, Setting 5: r=0.25r=-0.25).
+None of these correlations are statistically significant ($p>0.05$ for all settings), indicating universe captures complementary semantic cues.
+These results demonstrate that FVD and universe provide complementary evaluation signals: FVD captures distribution-level perceptual quality and temporal coherence, while universe quantifies task-specific semantic alignment through human-grounded assessment.
+H.2 Correlation with VBench Metrics
+Figure 17: Correlation between universe and VBench evaluation suite across eight environments and six dimensions.
+The human action dimension is omitted as it was constant and non-informative.
+Correlations remain consistently low to moderate (|r| < 0.4), with none of the correlations are statistically significant (all $p>0.05$).
+To validate that universe captures evaluative dimensions beyond existing automated benchmarks, we compare its human-aligned evaluation scores against VBench Huang et al.
+(2024), a comprehensive state-of-the-art benchmark for generative video evaluation.
+We computed correlations on identical rollout sets across eight experimental environments using six evaluation configurations, parameterized by task (Action Recognition or Character Recognition) and prompt type (binary, multiple-choice, open-ended).
+The analysis spans seven VBench dimensions: Temporal Flickering, Temporal Style, Background Consistency, Overall Consistency, Subject Consistency, Dynamic Degree, and Human Action.
+The Human Action dimension is omitted from visualization as it exhibited no variance across samples.
+Results.
+Figure 17 shows that correlations between universe and VBench metrics remain consistently low to moderate ($|r|<0.4$) across all valid comparisons.
+Average correlations range from -0.037 for Subject Consistency to 0.06 for Temporal Flickering, with most values falling within 0.1-0.4.
+Notably, the sign variability across settings (including both positive and negative correlations) indicates that universe captures complementary semantic cues not reflected in VBench’s quality dimensions.
+This divergence is expected: while VBench focuses on low-level perceptual quality and temporal coherence, universe explicitly targets semantic alignment.
+Summary.
+The consistent orthogonality observed across both analyses demonstrates that universe provides a complementary evaluation dimension to existing benchmarks.
+While standard distributional metrics (FVD) capture visual fidelity and text-to-video benchmarks (VBench) assess perceptual quality, universe quantifies action-conditioned semantic alignment through human-grounded evaluation.
+Comprehensive world model assessment therefore benefits from combining these approaches: existing metrics for perceptual coherence and universe for task-relevant semantic fidelity.